### v3

# Kernel Single Proxy Control for Deterministic Confounding

Abstract
We consider the problem of causal effect estimation with an unobserved confounder, where we observe a proxy variable that is associated with the confounder. Although Proxy causal learning (PCL) [4, 14] uses two proxy variables to recover the true causal effect, we show that a single proxy variable is sufficient for causal estimation if the outcome is generated deterministically, generalizing Control Outcome Calibration Approach (COCA) [23, 24]. We propose two kernel-based methods for this setting: the first based on the two-stage regression approach, and the second based on a maximum moment restriction approach. We prove that both approaches can consistently estimate the causal effect, and we empirically demonstrate that we can successfully recover the causal effect on challenging synthetic benchmarks.
1 Introduction
Causal learning involves estimating the impact of our actions on the world. For instance, we might want to understand how flight ticket prices affect sales, or how a specific medication increases the chance of recovery. The action we take is referred to as a “treatment,” which leads to a specific “outcome.” Determining the causal effect solely from observational data is often challenging, however, because the observed relationship between treatment and outcome can be influenced by an unobserved factor known as a “confounder.” In the case of the medical example, patient health might act as a confounder: if the patient is weaker, then he or she might require a higher dose of medicine, and yet have a lower chance of survival. Thus, from observing historical data, the dose may be negatively correlated with the chance of recovery; an instance of Simpson’s paradox. Hence, it is necessary to address the bias caused by the confounder to assess the true effect.
Although many methods for causal inference methods assume that there are no unmeasured confounders, this can be too restrictive, and it is often difficult to determine how the confounder affects treatment assignments and outcomes. A milder assumption is that we have access to proxy variables, which are related to the confounder in a specific way. In the proxy causal learning (PCL) framework [4, 14], it is shown that with two types of proxy variable, we can recover the true causal effect. The original work [14] provided a closed-form solution in the case of categorical treatments and outcomes, and derived a Fredholm equation of the first kind to characterize the continous case. Solutions for the latter case have been obtained where the causal relationships are modeled as nonlinear basis functions [4], RKHS functions [13, 20], and neural networks [25, 9, 7].
Another approach to using proxy variables is negative outcome control, which only requires one proxy variable, which is correlated with the confounder but not causally affected by the treatment. Although this requires less information than the PCL framework, we need to place stricter assumptions on the confounding structure. The most well-known negative outcome control is difference-in-difference (DiD) method [2, 3], which uses the pre-treatment outcome as the proxy and makes the parallel trends assumption, i.e. that the average trends in untreated potential outcomes for treatment and untreated groups are parallel. Although the DiD method is widely used, the parallel trends assumption may not be satisfied, since there might be an unobserved factor that affects both the trend and the treatment assignment.
To overcome this limitation, Tchetgen Tchetgen [23] proposed the Control Outcome Calibration Approach (COCA), which assumes the confounder is the potential outcome [19]. The method originally assumes a specific parametric form, known as rank preservation, which is relaxed in the recent Single Proxy Control work[24]. The latter approach still assumes a binary treatment, however, and that the confounder is the potential outcome.
Motivated by this, we propose Kernel Single Proxy Control, which can handle continuous treatments and general confounders. Our key assumption is deterministic confounding, which means that the outcome is deterministically generated given the treatment and the confounder. Note that this setting includes the original single proxy control, as the relation between the potential outcome and the observed outcome is deterministic. We show that the causal relation can be recovered by solving a similar functional equation as in single proxy control, which we propose to solve using a technique adapted from kernel PCL methods [13, 20]. We also derive a novel closed-form solution for PCL with improved numerical stability, which may be of independent interest.
This paper is structured as follows. In Section 2, we present an overview of the problem setting, and of relevant background material on kernel methods. In Section 3, we establish conditions under which the causal effect can be recovered from the observed variables. Then, in Section 4, we introduce our estimation procedures: namely, the two-stage least squares and moment matching approaches, as in Mastouri and Zhu et al. [13], with consistency results for each. In Section 5, we report the empirical performance of the proposed method on synthetic causal data. We also explore the case where the deterministic confounding assumption is violated, and show that our proposed methods can still obtain a reasonable estimate of the causal relationship.
2 Preliminary
Notation.
Throughout the paper, a capital letter (e.g. $A$) denotes a random variable, and we denote the set where a random variable takes values by the corresponding calligraphic letter (e.g. $\mathcal{A}$). The symbol $\mathbb{P}\left[\cdot\right]$ denotes the probability distribution of a random variable (e.g. $\mathbb{P}\left[A\right]$). We use a lowercase letter to denote a realization of a random variable (e.g. $a$). We denote the expectation over a random variable as $\mathbb{E}[\cdot]$ and $\|f\|_{\mathbb{P}\left[\cdot\right]}$ as the $L^{2}$-norm of a function $f$ with respect to $\mathbb{P}\left[\cdot\right]$; i.e. $\|f\|_{\mathbb{P}\left[A\right]}=\sqrt{\mathbb{E}_{A}\left[f^{2}(A)\right]}$.
2.1 Problem Settings
We begin by describing the problem setting. We observe the treatment $A\in\mathcal{A}$ and corresponding outcome $Y\in\mathcal{Y}$. We assume that there exists an unobserved confounder $U\in\mathcal{U}$ that affects both $A$ and $Y$. Our goal is to estimate the structural function $f_{\mathrm{struct}}$ defined as
$f_{\mathrm{struct}}(\tilde{a})=\mathbb{E}_{U}\left[\mathbb{E}_{Y}\left[Y|A=\tilde{a},U\right]\right],$
for a given test point $\tilde{a}\in\mathcal{A}$. This function is also known as the Dose-Response Curve or Average Potential Outcome. The challenge of estimating $f_{\mathrm{struct}}$ is that the confounder $U$ is not observable — we cannot estimate the structural function from observations $A$ and $Y$ alone. To deal with this, we assume access to the proxy $W$, which satisfies
Assumption 2.1.
We assume $W\mathop{\perp\!\!\!\perp}(A,Y)|U$.
Assumption 2.2.
For any square integrable function $l:\mathcal{U}\to\mathbb{R}$, the following conditions hold for all $a\in\mathcal{A}$:
$\mathbb{E}\left[l(U)\mid A=a,W=w\right]=0\quad\forall w\in\mathcal{W}$
$\quad\Leftrightarrow\quad l(u)=0\quad\mathbb{P}\left[U\right]\text{-}\mathrm{a.e.}$
$U$$A$$Y$$W$
Figure 1: Typical causal graph for the single proxy variable. Here, the bidirectional arrows mean that we allow an arrow in either direction, or even a common ancestor variable; and the thick arrow indicates a deterministic relationship between the variables.
Figure 1 shows the causal graph describing these relationships. We may additionally consider observable confounder, which is discussed in Appendix A.
The assumptions on the proxy $W$ are similar to those made in Proxy Causal Learning (PCL) [4, 14, 13, 20, 25, 7]. Our setting only assumes access to a single proxy $W$, however, while PCL requires an additional treatment-proxy $Z$ that satisfies $Z\mathop{\perp\!\!\!\perp}Y|A,U$. Instead of observing this proxy, we place deterministic confounding assumptions on the outcome $Y$, as follows.
Assumption 2.3.
There exists a deterministic function $\gamma_{0}(a,u)$ such that $Y=\gamma_{0}(A,U)$ and $\|\gamma_{0}(a,U)\|_{\mathbb{P}\left[U\right]}\leq\infty$ for all $a\in\mathcal{A}$.
This deterministic confounding assumption is unique in our setting: we will further discuss its necessity in Section 3. Note that it holds in many interesting cases. One is when $U$ represents the vector of the potential outcomes $U=\{Y^{(0)},\dots,Y^{(N_{A})}\}$ [19], where $\mathcal{A}=\{0,1,\cdots,N_{A}\}$. Then $\gamma_{0}(a,u)=\sum_{i=0}^{N_{A}}\mathbbm{1}(a=i)Y^{(i)}$ is deterministic given treatment $A$ and confounder $U$. This setting is studied in Control Outcome Calibration Approach (COCA) [24] when $N_{A}=2$. Our methods can be viewed as the generalization of the to continuous treatments and more general confounders $U$.
In COCA, the following additional assumption on the proxy $W$ is employed.
Assumption 2.4 ( 24, Assumption 2(iii)).
For test point $\tilde{a}\in\mathcal{A}$, we have $W\mathop{\perp\!\!\!\perp}A|Y^{(\tilde{a})}$.
This ensures that the proxy $W$ only affects the treatment through the confounder $U$, which can be a mechanism captured by the potential output to be estimated $Y^{(\tilde{a})}$. In Tchetgen Tchetgen et al. [24], it is discussed that municipality-level birth rates before the Zika virus outbreak satisfy the assumption when the treatment $A$ is the Zika epidemic status and the outcome $Y$ is the post-epidemic birth rate.
Instead of Assumption 2.4, we make a different assumption, which is similar to “informative treatment-proxy” assumption in PCL.
Assumption 2.5.
For any square integrable function $l:\mathcal{U}\to\mathbb{R}$, the following conditions hold for test point $\tilde{a}\in\mathcal{A}$:
$\mathbb{E}\left[l(U)\mid A=\tilde{a},Y=y\right]=0\quad\forall y\in\mathcal{Y}$
$\quad\Leftrightarrow\quad l(u)=0\quad\mathbb{P}\left[U\right]\text{-}\mathrm{a.e.}$
This gives an alternative understanding of our setting: We may use the outcome $Y$ as the “treatment proxy” $Z$ in classical PCL under our assumptions (notably deterministic confounding). See Appendix B for details. This assumption implies that the outcome can reveal the information of hidden confounder. One typical example of this is a time-evolving deterministic process. For example, in a chemical synthesis plant, the amount of the final product $Y$ can be computed given the perfect information of the plant state $U$ and the control signal $A$. These dynamics are governed by the chemical process, and thus, they are deterministic and invertible. Therefore, we can estimate the causal quantities in such cases given the proxy $W$, which reflects the state of the plant $U$. A related setting is (nonlinear) independent component analysis (ICA) [ICA], in which the latent has the same dimension as the observation, and $W$ serves a related function to the "additionally observed variable" in nonlinear ICA [khemakhem20nICA]. For this case, it is further assumed that the latent distribution factorizes into a product of marginals.
2.2 Reproducing Kernel Hilbert Space
For any space $\mathcal{F}\in\{\mathcal{A},\mathcal{Y},\mathcal{W}\}$, let $k_{\mathcal{F}}:\mathcal{F}\times\mathcal{F}\to\mathbb{R}$ be a positive semidefinite kernel. Here, we assume all kernel functions are characteristic [22] and bounded, $\max_{f,f^{\prime}\in\mathcal{F}}k_{\mathcal{F}}(f,f^{\prime})\leq\kappa$. We denote corresponding RKHS as $\mathcal{H}_{\mathcal{F}}$ and canonical feature as $\phi_{\mathcal{F}}$. The inner product and the norm in $\mathcal{H}_{\mathcal{F}}$ is denoted as $\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}_{\mathcal{F}}}$ and $\|\cdot\|_{\mathcal{H}_{\mathcal{F}}}$, respectively. We also define the tensor product as $\otimes$, which induces the Hilbert-Schmidt operator $f\otimes g:\mathcal{H}_{\mathcal{F}_{2}}\to\mathcal{H}_{\mathcal{F}_{1}}$ for $f\in\mathcal{H}_{\mathcal{F}_{1}},g\in\mathcal{H}_{\mathcal{F}_{2}}$, such that $(f\otimes g)h=\left\langle g,h\right\rangle_{\mathcal{H}_{\mathcal{F}_{2}}}f$ for $h\in\mathcal{H}_{\mathcal{F}_{2}}$. We denote the space of Hilbert-Schmidt operators of $\mathcal{H}_{\mathcal{F}_{2}}\to\mathcal{H}_{\mathcal{F}_{1}}$ as $\mathcal{H}_{\mathcal{F}_{1}\mathcal{F}_{2}}$ with Hilbert-Schmidt norm $\|\cdot\|_{\mathcal{H}_{\mathcal{F}_{1}\mathcal{F}_{2}}}$. We introduce the notation of $\phi_{\mathcal{F}_{1}\mathcal{F}_{2}}(f_{1},f_{2})=\phi_{\mathcal{F}_{1}}(f_{1})\otimes\phi_{\mathcal{F}_{2}}(f_{2})$. Note that it can be shown that $\mathcal{H}_{\mathcal{F}_{1}\mathcal{F}_{2}}$ is isometrically isomorphic to the product space $\mathcal{H}_{\mathcal{F}_{1}}\times\mathcal{H}_{\mathcal{F}_{2}}$.
3 Identification
In this section, we show that the structural functions can be estimated from observable variables $(Y,A,W)$ given assumptions in Section 2.1. We first present the general form of the identification results, and then discuss the special case when $(Y,W,U)$ are categorical or follow structural linear equations. Finally, we discuss the case where the outcome involves bounded additive noise.
3.1 Identification in the General Case
To obtain the structural function, we consider the bridge function $h_{0}(a,w)$, which is defined as the solution of the following functional equation,
$\mathbb{E}\left[h_{0}(a,W)|A=a,Y=y\right]=y$ (1)
Although the existence of such bridge function is assumed in Tchetgen Tchetgen et al. [24], we give the concrete condition of the existence as follows.
Theorem 3.1.
Given Assumptions 2.1, 2.2 and 2.3 and Assumptions D.3, D.4 and D.3 presented in Section D.3, then there exists a bridge function $h_{0}(a,w)$ satisfying (1).
The proof generalizes PCL identification [14] to the case where the outcome $Y$ is also the treatment proxy. Details can be found in Section D.4. Given bridge function $h_{0}(a,w)$, we can recover the structural function by taking the partial average of the bridge function as follows.
Theorem 3.2.
Given the bridge function $h_{0}(a,w)$ satisfying (1), and requiring either Assumption 2.4 or Assumption 2.5, then
$f_{\mathrm{struct}}(\tilde{a})=\mathbb{E}_{W}\left[h_{0}(\tilde{a},W)\right].$
The proof can be found in Section D.4. Note that we do not need both Assumption 2.4 and Assumption 2.5; either one is sufficient for identification. This analysis extends the results in Tchetgen Tchetgen et al. [24], who consider the case where Assumption 2.4 holds and the treatment $A$ is binary, $\mathcal{A}=\{0,1\}$. By contrast, in our analysis, we do not assume any specific form of $\gamma_{0}(a,u)$, which can be applied to a wider range of problems.
3.2 Identification in the Special Cases
To illustrate our method and the implications of our assumptions, we discuss the case where all variables are categorical and where the variables follow structural linear equations.
Discrete Variables
Assume each spaces $\mathcal{F}\in\{\mathcal{Y},\mathcal{W},\mathcal{U}\}$ are =$\mathcal{F}=\{1,\dots,d_{\mathcal{F}}\}$, and let $P(Y|\tilde{a},U),P(U)$ be the probability matrix $P(Y|\tilde{a},U)=(\mathbb{P}\left[Y=i|\tilde{a},U=j\right])_{ij}\in\mathbb{R}^{d_{\mathcal{Y}}\times d_{\mathcal{U}}},P(U)=(\mathbb{P}\left[U=i\right])_{i}\in\mathbb{R}^{d_{\mathcal{U}}}$. From the definition of the structural function, we have
$f_{\mathrm{struct}}(\tilde{a})=\bm{d}^{\top}P(Y|\tilde{a},U)P(U).$
where $\bm{d}=[1,2,\dots,d_{\mathcal{Y}}]^{\top}$. We cannot, however, directly compute this since $U$ is not observed. Instead, we aim to obtain the weights matrix $H_{a}$ that satisfies
$P(Y|\tilde{a},U)=H_{\tilde{a}}P(W|U),$ (2)
where $P(W|U)=(\mathbb{P}\left[W=i|U=j\right])_{ij}$ is the probability matrix of $\mathbb{P}\left[W|U\right]$. Given such a matrix $H_{a}$, the structural function can be written as
$f_{\mathrm{struct}}(\tilde{a})=\bm{d}^{\top}H_{\tilde{a}}P(W|U)P(U)=\bm{d}^{\top}H_{\tilde{a}}P(W),$
where $P(W)$ is a probability vector $P(W)=(\mathbb{P}\left[W=i\right])_{i}$, which can be estimated from observed proxy $W$.
We cannot directly solve (2) since it contains unobserved confounder $U$. Instead, we multiply both sides of (2) by another probability matrix $P(U|\tilde{a},Y)=(\mathbb{P}\left[U=i|\tilde{a},Y=j\right])_{ij}$ and consider
$P(Y|\tilde{a},U)P(U|\tilde{a},Y)$ $=H_{\tilde{a}}P(W|U)P(U|\tilde{a},Y)$
$=H_{\tilde{a}}P(W|\tilde{a},Y),$ (3)
where $P(W|\tilde{a},Y)=(\mathbb{P}\left[W=i|\tilde{a},Y=j\right])_{ij}$, and may be estimated from observable samples $(A,Y,W)$. The last equality holds from the assumption $W\mathop{\perp\!\!\!\perp}Y|A,U$. Note that if matrix $P(U|\tilde{a},Y)$ is right invertible (which is assumed in Assumption 2.5), the solution of (3) also satisfies (2).
The left-hand side of (3) still poses a problem, as we would need to observe $U$ in order to obtain $P(Y|\tilde{a},U)$ and $P(U|\tilde{a},Y)$. We can show, however, that the product of these two matrices is the identity matrix when the outcome $Y$ is deterministic given confounder $U$ and treatment $A$.
Lemma 3.3.
Under Assumption 2.3, we have
$P(Y|\tilde{a},U)P(U|\tilde{a},Y)=I_{d_{\mathcal{Y}}}$
where $I_{d_{\mathcal{Y}}}\in\mathbb{R}^{d_{\mathcal{Y}}\times d_{\mathcal{Y}}}$ is the identity matrix.
The proof is presented in Section D.1. This is the reason that we need deterministic confounding; otherwise, we cannot solve (3) from observable variables $(A,Y,W)$. From 3.3 and (3), we can see that $H_{\tilde{a}}$ is the left-inverse matrix of $P(W|\tilde{a},Y)$. In other words, the $(i,j)$-element in $H_{\tilde{a}}$, denoted as $(H_{\tilde{a}})_{ij}$, satisfies
$\delta_{yi}=\sum_{j=1}^{d_{\mathcal{W}}}(H_{\tilde{a}})_{ij}\mathbb{P}(W=j|A=\tilde{a},Y=i)\quad\forall y\in\mathcal{Y}$
for Kronecker delta $\delta_{ij}$, which coincides with the bridge function (1).
Linear Relationships
Assume all variables are in $\mathbb{R}$ and follow the linear structural equations:
$U=\varepsilon_{u},\,W=\alpha_{wu}U+\varepsilon_{w},$
$A=\alpha_{xu}U+\varepsilon_{a},\,Y=\alpha_{ya}A+\alpha_{yu}U+\varepsilon_{y},$
where $\varepsilon_{u},\varepsilon_{w},\varepsilon_{a},\varepsilon_{y}$ are independent random variables with zero mean. This setting is discussed in Kuroki2014Mesurement, and it shows we cannot recover $\alpha_{ya}$ unless we can estimate $\alpha^{2}_{wu}\mathrm{Var}[U]$. Our work avoids this requirement, however, by making the deterministic assumption. If we consider a linear bridge function $h_{0}(a,w)=h_{a}a+h_{w}w$, we have
$\mathbb{E}\left[h_{0}(a,W)|A=a,Y=y\right]$
$\quad=h_{a}a+h_{w}\mathbb{E}\left[W|A=a,Y=y\right]$
$\quad=h_{a}a+h_{w}\alpha_{wu}\mathbb{E}\left[U|A=a,Y=y\right]$
$\quad=h_{a}a+h_{w}\alpha_{wu}(y-\alpha_{ya}a)/\alpha_{yu},$
where the last equation holds if and only if $\varepsilon_{y}=0$, which corresponds to deterministic confounding Assumption 2.3. Hence, if we solve (1), we have
$h_{a}=\alpha_{ya},\quad h_{w}=\alpha_{yu}/\alpha_{wu},$
when $\alpha_{wu}\neq 0$ and $\alpha_{yu}\neq 0$, which correspond to Assumption 2.2 and Assumption 2.5, respectively. From this, we show that we can obtain the effect $\alpha_{ya}$ as the coefficient of the bridge function $h_{a}$ if the outcome is given deterministically.
3.3 Analysis of Stochastic Confounding
One natural question is whether we can recover the structural function when the outcome is given stochastically. Here, we discuss when the outcome involves bounded additive noise.
Assumption 3.4 (Bounded Additive Noise).
Let $\gamma_{0}$ be $\gamma_{0}(a,u)=\mathbb{E}\left[Y|a,u\right]$ where $\|\gamma_{0}(a,U)\|_{\mathbb{P}\left[U\right]}\leq\infty$ for all $a\in\mathcal{A}$. Assume that outcome is given as $Y=\gamma_{0}(A,U)+\varepsilon$ where $\mathbb{E}\left[\varepsilon\right]=0,\,|\varepsilon|\leq M$.
In this case, we cannot recover the true structural function by taking the partial average of the bridge function. Instead, we provide the upper-bound of the estimation error as follows.
Theorem 3.5.
Assume Assumptions 2.1 and 3.4, and assume that there exists $\Xi\leq\infty$ for all bounded function $\ell(u)$,
$\sup_{u}|\ell(u)|\leq\Xi\sup_{y\in\mathcal{Y}}{\mathbb{E}\left[\ell(U)|A=a,Y=y\right]}$ (4)
for all $(a,y)\in\mathcal{A}\times\mathcal{Y}$. If bridge function $h_{0}$ satisfying (1) exists, for all $a\in\mathcal{A}$, we have
$|f_{\mathrm{struct}}(a)-\mathbb{E}_{W}\left[h(a,W)\right]|\leq M\Xi$
The proof can be found in Section D.5. Note that the condition in (4) is the sufficient condition for Assumption 2.5. One important difference compared to 3.2 is that the existence of bridge function $h_{0}(a,w)$ is not guaranteed in 3.5. This is because we need Assumption 2.3 to prove the existence of the bridge function.
Although 3.5 only upper-bounds the error, the experiments in Section 5 show promising empirical results. Further theoretical analysis on this stochastic case is left as future work.
4 Methods
In this section, we present our proposed methods to estimate the bridge function. We introduce two algorithms, Single Kernel Proxy Variable (SKPV) and Single Proxy Maximum Moment Restriction (SPMMR), which estimate the bridge function with different approaches.
From 3.2, we can see that it is sufficient to estimate the bridge function to obtain the structural function, which can be done by solving functional equation (1) using samples $\{a_{i},w_{i},y_{i}\}_{i=1}^{n}$. This problem is ill-posed when we consider a sufficiently rich functional space (see discussion in Nashed and Wahba [16]). Here, we assume that the bridge function $h$ is in RKHS $\mathcal{H}_{\mathcal{A}\mathcal{W}}$ and add a smoothness requirement. We will discuss the case where the bridge function $h$ can be represented as neural network in Appendix C.
We propose two approaches to solve (1): A Two-stage regression approach [20, 25, 13], which minimizes the $L^{2}$-distance between both sides of (1); and a Maximum Moment Restriction approach [13, 9, 7], which minimizes the maximum moment of the deviation in (1). We also derive novel closed-form solutions in these approaches, which improve the numerical stability compared to the existing solutions.
4.1 Two-Stage Regression Approach
In two-stage regression, we solve (1) by minimizing the loss
$\mathcal{L}^{\mathrm{2SR}}(h)=\mathbb{E}_{AY}\left[(Y\!-\!\mathbb{E}\left[h(A,W)|A,Y\right])^{2}\right]\!+\!\eta\|h\|^{2}_{\mathcal{H}_{\mathcal{A}\mathcal{W}}},$
where $\eta\|h\|^{2}_{\mathcal{H}_{\mathcal{A}\mathcal{W}}}$ is the regularizing term. This loss cannot be directly minimized as it involves the conditional expectation $\mathbb{E}\left[h(A,W)|A,Y\right]$. To deal with this, Mastouri and Zhu et al. [13]and Singh [20] employed a two-stage regression approach, which first estimates conditional expectation, and then minimizes the loss. We apply this approach in our single proxy setting, and propose the Single Kernel Proxy Variable (SKPV) method. Specifically, if $h\in\mathcal{H}_{\mathcal{A}\mathcal{W}}$, we have
$\mathbb{E}\left[h(a,W)|A=a,Y=y\right]$
$\quad=\left\langle h,\phi_{\mathcal{A}}(a)\otimes\mu_{W|A,Y}(a,y)\right\rangle_{\mathcal{H}_{AW}}.$
Here, $\mu_{W|A,Y}(a,y)$ is known as a conditional mean embedding [21, 6, 11, 17], defined as
$\mu_{W|A,Y}(a,y)=\mathbb{E}\left[\phi_{\mathcal{W}}(W)|A=a,Y=y\right]$
Under the regularity condition that $\mathbb{E}\left[g(W)|A=a,Y=y\right]$ is in $\mathcal{H}_{\mathcal{A}\mathcal{Y}}$ for all $g\in\mathcal{H}_{\mathcal{W}}$, there exists an operator $C_{W|A,Y}$ such that $\mu_{W|A,Y}=C_{W|A,Y}(\phi_{\mathcal{A}}(a)\otimes\phi_{\mathcal{Y}}(y))$. This operator can be estimated from the samples $\{w_{i},a_{i},y_{i}\}_{i=1}^{n}$ [21, 6, 11, 17] by minimizing the loss $\hat{C}_{W|A,Y}=\operatorname*{arg\,min}_{C\in\mathcal{H}_{\mathcal{W}(\mathcal{A}\mathcal{Y})}}{\hat{\mathcal{L}}}^{\mathrm{cond}}(C)$ defined as
^${\hat{\mathcal{L}}}^{\mathrm{cond}}(C)$
$~{}=\frac{1}{n}\sum_{i=1}^{n}\|\phi_{\mathcal{W}}(w_{i})-C\phi_{\mathcal{A}\mathcal{Y}}(a_{i},y_{i})\|_{\mathcal{H}_{\mathcal{W}}}^{2}\!\!+\lambda\|C\|^{2}_{\mathcal{H}_{\mathcal{W}(\mathcal{A}\mathcal{Y})}},$
where $\phi_{\mathcal{A}\mathcal{Y}}(a,y)=\phi_{\mathcal{A}}(a)\otimes\phi_{\mathcal{Y}}(y)$ is the tensor feature, $\lambda>0$ is the regularizing coefficient, $\mathcal{H}_{\mathcal{W}(\mathcal{A}\mathcal{Y})}$ denotes the space of Hilbert-Schmidt operators that maps $\mathcal{H}_{\mathcal{A}\mathcal{Y}}$ to $\mathcal{H}_{\mathcal{W}}$, and $\|\cdot\|_{\mathcal{H}_{\mathcal{W}(\mathcal{A}\mathcal{Y})}}$ is the norm of that space. This is known as stage 1 regression as we regress the feature $\phi_{\mathcal{W}}(w_{i})$ on tensor product features $\phi_{\mathcal{A}}(a_{i})\otimes\phi_{\mathcal{Y}}(y_{i})$. The closed-from solution is
$\hat{C}_{W|A,Y}=\hat{C}_{W,(A,Y)}(\hat{C}_{(A,Y),(A,Y)}+\lambda I)^{-1}$
where
$\hat{C}_{W,(A,Y)}$ $=\frac{1}{n}\sum_{i=1}^{n}\phi_{\mathcal{W}}(w_{i})\otimes\phi_{\mathcal{A}\mathcal{Y}}(a_{i},y_{i})$
$\hat{C}_{(A,Y),(A,Y)}$ $=\frac{1}{n}\sum_{i=1}^{n}\phi_{\mathcal{A}\mathcal{Y}}(a_{i},y_{i})\otimes\phi_{\mathcal{A}\mathcal{Y}}(a_{i},y_{i}).$
Given this, we minimize $\mathcal{L}^{\mathrm{2SR}}$ using the estimate in stage 1 regression. Specifically, we obtain the bridge function estimate $\hat{h}$ as $\hat{h}=\operatorname*{arg\,min}_{h\in\mathcal{H}_{\mathcal{A}\mathcal{W}}}{\hat{\mathcal{L}}}^{\mathrm{2SR}}(h)$ given stage 2 data $\{\dot{a}_{i},\dot{y}_{i}\}_{i=1}^{m}$, where the loss is defined as
^${\hat{\mathcal{L}}}^{\mathrm{2SR}}(h)\!=\!\frac{1}{n}\sum_{i=1}^{n}\left(\dot{y}_{i}\!-\!\left\langle h,\hat{\phi}_{\mathrm{2}}(\dot{a}_{i},\dot{y}_{i})\right\rangle_{\mathcal{H}_{\mathcal{A}\mathcal{W}}}\right)^{2}\!+\!\eta\|h\|^{2}_{\mathcal{H}_{\mathcal{A}\mathcal{W}}},$
where $\hat{\mu}_{W|A,Y}(a,y)=\hat{C}_{W|A,Y}\phi_{\mathcal{A}\mathcal{Y}}(a,y)$ is the estimated conditional mean embedding and $\hat{\phi}_{\mathrm{2}}(\dot{a}_{i},\dot{y}_{i})=\phi_{\mathcal{A}}(\dot{a}_{i})\otimes\hat{\mu}_{W|A,Y}(\dot{a}_{i},\dot{y}_{i})$ is the stage 2 tensor feature.
This regression is known as stage 2 regression. We may use the same samples in stage 1 and 2 regressions, but it is reported that splitting samples between stage 1 and 2 regression can reduce bias in the estimation [1]. Furthermore, since we do not need samples of $W$ in stage 2 regression, we may use an additional source of data for the samples $\{\dot{a}_{i},\dot{y}_{i}\}$. We can obtain the closed-form solution of two-stage regression as follows:
Theorem 4.1.
Given stage 1 samples $\{w_{i},a_{i},y_{i}\}_{i=1}^{n}$ and stage 2 samples $\{\dot{a}_{i},\dot{y}_{i}\}_{i=1}^{m}$, and regularizing parameter $(\lambda,\eta)$, the minimizer of ^${\hat{\mathcal{L}}}^{\mathrm{2SR}}$ is given as
$\hat{h}(a,w)=\bm{\alpha}^{\top}\bm{k}(a,w),\quad\bm{\alpha}=(M+m\eta I)^{-1}\dot{\bm{y}}$
where $\dot{\bm{y}}=(\dot{y}_{1},\dots,\dot{y}_{m})^{\top}\in\mathbb{R}^{m}$, and $M\in\mathbb{R}^{m\times m}$ and $\bm{k}(a,w)\in\mathbb{R}^{m}$ are defined as
$M=K_{\dot{A}\dot{A}}\odot(B^{\top}K_{WW}B),$
$\bm{k}(a,w)=\bm{k}_{\dot{A}}(a)\odot(B^{\top}\bm{k}_{W}(w)),$
$B=(K_{AA}\odot K_{YY}+n\lambda I)^{-1}(K_{A{\dot{A}}}\odot K_{Y\dot{Y}}).$
Here, for $F\in\{A,W,Y\}$, we denote the stage 1 kernel matrix as $K_{FF}=(k_{\mathcal{F}}(f_{i},f_{j}))_{ij}\in\mathbb{R}^{n\times n}$ where $k_{\mathcal{F}}$ and $f_{i}$ are the corresponding space and stage 1 samples. Similarly, for $\dot{F}\in\{\dot{A},\dot{Y}\}$, we denote the stage 2 kernel matrix as $K_{\dot{F}\dot{F}}=(k_{\mathcal{F}}(\dot{f}_{i},\dot{f}_{j}))_{ij}\in\mathbb{R}^{n\times n}$ where $k_{\mathcal{F}}$ and $f_{i}$ are the corresponding space and stage 2 samples. We further denote $K_{F\dot{F}}=(k_{\mathcal{F}}(f_{i},\dot{f}_{j}))_{ij}\in\mathbb{R}^{n\times m}$ and $\bm{k}_{\dot{A}}(a)=(k(\dot{a}_{i},a))_{i}\in\mathbb{R}^{m},\bm{k}_{W}(w)=(k(w_{i},w))_{i}\in\mathbb{R}^{n}$
The proof is in Appendix E, where we further show that this solution is equivalent to the result of Mastouri and Zhu et al. [13, Proposition 2], which involves learning $nm$ parameters, while ours only requires learning $n$ parameters. Our solution is moreover slightly different from Singh [20, Algorithm 1], which takes the form $\bm{\alpha}=(MM+m\eta M)^{-1}M\dot{\bm{y}}$. Although these solutions co-incide when $M$ is invertible, we claim that our solution is more numerically stable, especially when $M$ is close to singular. The computation requires $O(n^{3}+m^{3})$, which can be reduced via the usual Cholesky or Nystrom techniques.
Given estimated bridge function $\hat{h}(\tilde{a},w)$, we can estimate the structural function as $\hat{f}_{\mathrm{struct}}(\tilde{a})=\frac{1}{n}\sum_{i=1}^{n}\hat{h}(\tilde{a},w_{i}).$
4.2 Maximum Moment Restriction Approach
Another approach to solving (1) is to consider moment restriction, which minimizes the maximum moment of $\mathbb{P}\left[Y-h(a,W)|A,Y\right]$. We apply this approach to our setting and propose Single Proxy Maximum Moment Restriction (SPMMR). Specifically, we consider the following minimax problem.
$\operatorname*{arg\,min}_{h\in\mathcal{H}_{\mathcal{A}\mathcal{W}}}\max_{\begin{subarray}{c}g\in\mathcal{H}_{\mathcal{A}\mathcal{Y}},\\ \|g\|_{\mathcal{H}_{\mathcal{A}\mathcal{Y}}}\leq 1\end{subarray}}(\mathbb{E}\left[(Y-h(A,W))g(A,Y)\right])^{2}.$
As shown by Mastouri and Zhu et al. [13, Lemma 1], the unique solution to this minimax problem is the solution of (1). Furthermore, using a similar discussion to Mastouri and Zhu et al. [13, Lemma 2], we can show that the maximum over $g$ can be computed in closed form as
$\mathcal{L}^{\mathrm{MMR}}(h)$
$=\max_{g\in\mathcal{H}_{\mathcal{A}\mathcal{Y}},\|g\|_{\mathcal{H}_{\mathcal{A}\mathcal{Y}}}\leq 1}(\mathbb{E}\left[(Y-h(A,W))g(A,Y)\right])^{2}$
$=\!\mathbb{E}\left[\Delta_{Y,A,W}(h)\Delta_{Y^{\prime},A^{\prime},W^{\prime}}(h)k_{\mathcal{A}}(A,A^{\prime})k_{\mathcal{Y}}(Y,Y^{\prime})\right],$
where $\Delta_{Y,A,W}(h)=(Y-h(A,W))$ and $(W^{\prime},A^{\prime},Y^{\prime})$ are independent copies of $(W,A,Y)$, respectively. Empirically, we minimize $\mathcal{L}^{\mathrm{MMR}}$ with regularization as $\hat{h}=\operatorname*{arg\,min}_{h\in\mathcal{H}_{\mathcal{A}\mathcal{W}}}{\hat{\mathcal{L}}}^{\mathrm{MMR}}$
^${\hat{\mathcal{L}}}^{\mathrm{MMR}}$ $=\sum_{i,j=1}^{n}\frac{\Delta_{i}\Delta_{j}}{n^{2}}k_{\mathcal{A}}(a_{i},a_{j})k_{\mathcal{Y}}(y_{i},y_{j})\!+\!\eta\|h\|^{2}_{\mathcal{H}_{\mathcal{A}\mathcal{W}}},$
where $\Delta_{i}=(y_{i}-h(a_{i},w_{i}))$. The closed-form solution of $\hat{h}$ can be obtained as follows.
Theorem 4.2.
Given samples $\{w_{i},a_{i},y_{i}\}_{i=1}^{n}$, the minimizer of ^${\hat{\mathcal{L}}}^{\mathrm{MMR}}$ is given as $\hat{h}(a,w)=\bm{\alpha}^{\top}\bm{k}(a,w)$, where weight $\bm{\alpha}\in\mathbb{R}^{n}$ is given as
$\bm{\alpha}=\sqrt{G}\left(\sqrt{G}L\sqrt{G}+n^{2}\eta I\right)^{-1}\sqrt{G}\bm{y}.$
Here, $\bm{y}=(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$ and
$L=K_{AA}\odot K_{WW},\quad G=K_{AA}\odot K_{YY},$
$\bm{k}(a,w)=(k_{\mathcal{A}}(a_{i},a)k_{\mathcal{W}}(w_{i},w))_{ij}\in\mathbb{R}^{n}$
with $K_{FF}=(k_{\mathcal{F}}(f_{i},f_{j}))_{ij}\in\mathbb{R}^{n\times n}$ for $F\in\{W,A,Y\}$ and corresponding kernel function $k_{\mathcal{F}}$ and samples $f_{i}$, and $\sqrt{G}$ denotes the square root of $G=\sqrt{G}\sqrt{G}$.
The derivation is given in Appendix E. Again, this slightly differs from the original solution in Mastouri and Zhu et al. [13], which is $\bm{\alpha}=(LGL+n^{2}\eta L)^{-1}LG\bm{y}$, but they are identical if $L,G$ are non-singular. We claim that our solution is more numerically stable when the condition number of $L$ is large. The computational complexity is $O(n^{3})$, which can be reduced via the usual Cholesky or Nystrom techniques. We can use the empirical average of the estimated bridge function to obtain the structural function as in SKPV.
4.3 Consistency
We derive the consistency in this section. The proof is mostly adopted from Singh [20] and Mastouri et al. [13].
Two-stage Regression
To derive the consistency of two-stage regression, we make the following assumptions.
Assumption 4.3.
Let operator $C_{\mu}$ be
$C_{\mu}=\mathbb{E}\left[\phi_{\mathrm{2}}(A,Y)\otimes\phi_{\mathrm{2}}(A,Y)\right]$
for $\phi_{\mathrm{2}}(A,Y)=\phi_{\mathcal{A}}(A)\otimes\mu_{W|A,Y}(A,Y)\in\mathcal{H}_{\mathcal{A}\mathcal{W}}$ and its eigen-decomposition be $\{\eta_{i},\phi_{i}\}$. We assume $\eta_{i}$, the $i$-th eigenvalue, can be bounded by $\eta_{i}\leq Ci^{-b_{0}}$ for constant $C,b_{0}$. Furthermore, we assume that there exists $g\in\mathcal{H}_{\mathcal{A}\mathcal{W}}$ such that $h_{0}=C_{\mu}^{\frac{c_{0}-1}{2}}g$ for some $c_{0}\in(1,2]$.
This assumption is made in the analysis of various kernel-based methods [20, 13, 5]. The constant $b_{0}$ is known as the effective dimension of $\mathcal{H}_{\mathcal{A}\mathcal{W}}$, and $b_{0}=\infty$ when RKHS $\mathcal{H}_{\mathcal{A}\mathcal{W}}$ is of finite dimension. The constant $c_{0}$ measures the smoothness of the bridge function: a larger $c_{0}$ means a smoother bridge function. We assume a similar condition on operator $C_{W|A,Y}$.
Assumption 4.4.
Let the covariance operator of $\mathcal{H}_{\mathcal{A}\mathcal{Y}}$ be $C_{(A,Y),(A,Y)}=\mathbb{E}\left[\phi_{\mathcal{A}\mathcal{Y}}(A,Y)\otimes\phi_{\mathcal{A}\mathcal{Y}}(A,Y)\right]$. We assume $\eta_{i}$, the $i$-th eigenvalue of $C_{(A,Y),(A,Y)}$, can be bounded by $\eta_{i}\leq Ci^{-b_{1}}$ for constant $C,b_{1}$. Furthermore, we assume that there exists $G\in\mathcal{H}_{\mathcal{W}(\mathcal{A}\mathcal{Y})}$ that $C_{W|A,Y}=G\circ C^{\frac{c_{1}-1}{2}}_{(A,Y),(A,Y)}$ for some $c_{1}\in(1,2]$.
Here, we denote $*{}^{*}$ as the adjoint of the operator. Given these assumptions, we obtain the following consistency result.
Proposition 4.5.
Assume Assumptions 4.3 and 4.4. Given stage 1 samples $\{w_{i},a_{i},y_{i}\}_{i=1}^{n}$ and stage 2 samples $\{\dot{a}_{i},\dot{y}_{i}\}_{i=1}^{m}$, and let $\lambda=n^{-\frac{1}{c_{0}+1/b_{0}}}$ and $n=m^{\xi\frac{c_{0}+1/b_{0}}{c_{0}-1}}$ for some $\xi>0$. Then,
$\sup_{a,w}|h_{0}(a,w)-\hat{h}(a,w)|=O_{P}(m^{-\frac{\xi c_{1}-1}{2(c_{1}+1)}}),$
with $\eta=m^{-\frac{\xi}{c_{1}+3}}$ when $\xi\leq\frac{c_{1}+3}{c_{1}+1/b_{1}}$. Otherwise, we have
$\sup_{a,w}|h_{0}(a,w)-\hat{h}(a,w)|=O_{P}(m^{-\frac{c_{1}-1}{2(c_{1}+1/b_{1})}}),$
with $\eta=m^{-\frac{1}{c_{1}+1/{b_{1}}}}$. These upper bounds can be translated to the estimation error of structural function as
$|f_{\mathrm{struct}}(\tilde{a})-\hat{f}_{\mathrm{struct}}(\tilde{a})|$
$\quad\leq\sup_{w}|h_{0}(\tilde{a},w)-\hat{h}(\tilde{a},w)|+O_{P}(n^{-\frac{1}{2}}),$
under the condition of 3.2.
The proof is a direct application of Theorem 3 in Singh [20]. From this, we can see that SKPV can consistently recover the bridge function and structural function. As in Singh [20], we observe that the best data ratio is $n=m^{\frac{c_{1}+3}{c_{1}+1/b_{1}}\frac{c_{0}+1/b_{0}}{c_{0}-1}}$ from which $n\gg m$. The best rate for bridge function estimation under this data size ratio is $O_{P}(m^{-\frac{c_{1}-1}{2(c_{1}+1/b_{1})}})$ which is the same as for kernel ridge regression [5].
Maximum Moment Restriction
For maximum moment restriction approach, we use the same assumption as Mastouri and Zhu et al. [13, Assumption 16].
Assumption 4.6.
Let $\{\eta_{i},\phi_{i},\psi_{i}\}$ be a singular decomposition of operator $T=\mathbb{E}\left[\phi_{\mathcal{A}\mathcal{W}}(A,W)\otimes\phi_{\mathcal{A}\mathcal{Y}}(A,Y)\right]$. We assume $h_{0}$ is in a $\gamma$-regularity space $\mathcal{H}^{\gamma}_{\mathcal{A}\mathcal{W}}$ that is defined as
$\mathcal{H}^{\gamma}_{\mathcal{A}\mathcal{W}}=\left\{f\in\mathcal{H}_{\mathcal{A}\mathcal{W}}~{}\middle|~{}\sum_{i=1}^{\infty}\frac{\left\langle f,\phi_{i}\right\rangle_{\mathcal{H}_{\mathcal{A}\mathcal{W}}}^{2}}{\eta_{i}^{2\gamma}}\right\}$
This assumption is similar to Assumption 4.3: a larger $\gamma$ means a smoother bridge function. Given this, we can show the consistency of the SPMMR method.
Proposition 4.7.
Given Assumption 4.6 and samples $\{w_{i},a_{i},y_{i}\}$ size of $n$. Then, we have
$\|\hat{h}-h_{0}\|_{\mathcal{H}_{\mathcal{A}\mathcal{W}}}\leq O\left(n^{-\frac{1}{2}\min\left(\frac{1}{2},\frac{\gamma}{\gamma+2}\right)}\right)$
with $\lambda=n^{-\frac{1}{2}\min\left(\frac{1}{2},\frac{2}{\gamma+2}\right)}$.
The proof is the direct application of Mastouri and Zhu et al. [13, Theorem 3]. From this, we can see that the best rate $O(n^{-1/4})$ is achieved when $\gamma\geq 2$. Again, this best rate is the same as kernel ridge regression or SKPV. Furthermore, Mastouri and Zhu et al. [13] discussed the relation of the two-stage regression approach and maximum moment restriction approach in the asymptotic case; this relation also applies to our SKPV and SPMMR.
5 Experiments
In this section, we present the empirical performance of our single proxy method in a synthetic setting. Firstly, we discuss hyperparameter tuning for SKPV and SPMMR. Then, we present the results for the synthetic setting.
5.1 Regularization Parameter Tuning
We follow the procedure in Mastouri and Zhu et al. [13]to tune regularization parameter $(\lambda,\eta)$ for SKPV and $\lambda$ in SPMMR.
Tuning $(\lambda,\eta)$ in SKPV
For selecting the best $\lambda$ in stage 1 regression, we use leave-one-out error for ^${\hat{\mathcal{L}}}^{\mathrm{cond}}$. This can be computed as the closed-form solution
^${\hat{\mathcal{L}}}^{\mathrm{cond-loo}}(\lambda)=\mathrm{tr}({\tilde{H}_{1}}^{-1}H_{1}K_{WW}H_{1}{\tilde{H}_{1}}^{-1})$
where
$H_{1}=I-(K_{AA}\odot K_{YY})(K_{AA}\odot K_{YY}+n\lambda I)^{-1},$
and $\tilde{H}_{1}$ is the diagonal matrix with the same diagonals $H_{1}$. Given the best parameter $\lambda=\operatorname*{arg\,min}{\hat{\mathcal{L}}}^{\mathrm{cond-loo}}$, we can tune the second stage regularizer $\eta$ by using the stage 1 data as “validation data” in stage 2 loss:
$\eta=\operatorname*{arg\,min}\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\left\langle\hat{h},\phi_{\mathrm{2}}(a_{i},y_{i})\right\rangle_{\mathcal{H}_{\mathcal{A}\mathcal{W}}}\right)^{2},$
where $\hat{h}$ is the solution in 4.1.
Tuning $\lambda$ in SPMMR
Since the empirical loss ^${\hat{\mathcal{L}}}^{\mathrm{MMR}}$ is a V-statistic, we cannot use leave-one-out error to tune $\lambda$ in SPMMR. We select the best $\lambda$ by minimizing the empirical loss ^${\hat{\mathcal{L}}}^{\mathrm{MMR}}$ over a held-out validation set.
5.2 Empirical Experiments
We consider the following data generation processes:
$U$ $\sim\mathrm{Unif}(-1,1),~{}~{}$ $A$ $=\Phi(U)+\varepsilon_{1}$
$Y$ $=\sin(2\pi U)+A^{2}-0.3,~{}~{}$ $W$ $=\exp(U)+\varepsilon_{2}$
Here, $\Phi$ is the Gaussian error function, and $\varepsilon_{1}\sim\mathcal{N}(0,(0.1)^{2}),\varepsilon_{2}\sim\mathcal{N}(0,(0.05)^{2})$. The true structural function is $f_{\mathrm{struct}}(a)=a^{2}-0.3$.
Table 1: Mean squared error and standard deviation for various data sizes.
Data size $1000$ $5000$
CEVAE 0.120(0.005) 0.309 (0.003)
SKPV 0.043(0.001) 0.041(0.001)
SPMMR 0.055(0.000) 0.055(0.000)
To estimate the structural function, we apply SKPV and SPMMR. We used the Gaussian kernel, where the bandwidth is tuned using the median trick. We selected the regularizers using the procedure described in Section 5.1. We split data evenly for stage 1 and 2 regression in SKPV. We compare proposed methods to CEVAE [12], which uses a VAE [8] to recover the distribution of confounder $U$ from the “proxy” $W$. Although CEVAE does not always provide consistent causal effect estimates [18], it has shown strong empirical performance in a number of settings.
The estimation loss is summarized in Table 1, in which SKPV and SPMMR consistently outperform CEVAE when the data size is set to 1000 and 5000. In our experiment, SKPV performs better than SPMMR.
Noisy Outcome
We further test performance in a case where the deterministic confounding assumption (Assumption 2.3) is violated. We add Gaussian noise $\mathcal{N}(0,\sigma^{2})$ to the outcome, where we vary the noise levels in $\sigma\in\{0.1,1.0\}$. We again generated 1000 samples and repeated each noise level for 10 times. The mean squared error in estimating the structural function is presented in Table 2. From this, we can see that SKPV and KPMMR can still correct confounding bias even if the outcome contains noise. Overall, we hypothesize that the effect of the additive error is minimal as suggested by 3.5.
Table 2: Mean squared error and standard deviation for various noise levels.
Noise level $\sigma=0.1$ $\sigma=1.0$
CEVAE 0.111(0.003) 0.124(0.006)
SKPV 0.047(0.001) 0.051(0.002)
SPMMR 0.055(0.001) 0.081(0.001)
High-Dimensional Proxy
We also consider the case where the proxy is high-dimensional. We used the same data generation process for $(A,Y,U)$ with deterministic confounding, and replace the proxy $W$ with MNIST images, where the digit label is chosen as $\lfloor 5U+5\rfloor$. Results with data of size 1000 and 5000 are summarized in Table 3.
Table 3: Mean squared error and standard deviation when the proxy is MNIST images.
Data size $1000$ $5000$
CEVAE 10.43(5.080) 72.22(40.64)
SKPV 0.124(0.003) 0.121(0.001)
SPMMR 0.218(0.006) 0.224(0.002)
Since the proxy $W$ is high-dimensional, all methods perform worse than in the original settings (Table 3). In particular, the performance of CEVAE is highly unstable, consistent with the observation in [18] that for high dimensional $W$, the autoencoder can prioritise quality of reconstruction of $W$ over causal effect estimation. SKPV and SPMMR perform considerably better than CEVAE in this setting.
6 Conclusion
In this paper, we propose single proxy causal learning, which only requires a single proxy variable. Our theoretical analysis shows that the bridge function can recover the true structural function if the outcome is generated in a deterministic way. Motivated by this, we propose two methods to estimate the bridge function, named SKPV and SPMMR, which are shown to converge to the true bridge function as the number of samples increases. We also empirically show that we can still apply these methods in the additive noise case, and demonstrate a bound on the performance penalty incurred through violation of the deterministic outcome assumption.
Appendix A Observable Confounder
$U$$A$$Y$$W$$X$
Figure 2: Typical causal graph for the single proxy variable with observable confounder $X$. Here, the bidirectional arrows mean that we allow an arrow in either direction, or even a common ancestor variable; and the thick arrow indicates a deterministic relationship between the variables.
In this appendix, we discuss the case where there exists the observable confounder $X$. The causal graph in this case is given in Figure 2. Now, the structural function is defined as
$f_{\mathrm{struct}}(\tilde{a})=\mathbb{E}_{X,U}\left[\mathbb{E}\left[Y|A=\tilde{a},X,U\right]\right].$
Given the $X\in\mathcal{X}$, we would solve the bridge function $h_{0}(a,x,w):\mathcal{A}\times\mathcal{X}\times\mathcal{W}\to\mathcal{Y}$,
$\mathbb{E}\left[h_{0}(a,x,W)|A=a,X=x,Y=y\right]=y\quad\forall a\in\mathcal{A},y\in\mathcal{Y},x\in\mathcal{X}$ (5)
Such bridge function $h_{0}$ exists when the proxy satisfies similar conditions.
Assumption A.1.
We assume $W\mathop{\perp\!\!\!\perp}(A,Y)|U,X$.
Assumption A.2.
For any square integrable function $l:\mathcal{U}\to\mathbb{R}$, the following conditions hold for all $a\in\mathcal{A}$ and $x\in\mathcal{X}$:
$\mathbb{E}\left[l(U)\mid A=a,W=w,X=x\right]=0\quad\forall w\in\mathcal{W}\quad\Leftrightarrow\quad l(u)=0\quad\mathbb{P}\left[U\right]\text{-}\mathrm{a.e.}$
Assumption A.3.
There exists a deterministic function $\gamma_{0}(a,x,u)$ such that $Y=\gamma_{0}(A,U,X)$ and $\|\gamma_{0}(a,x,U)\|_{\mathbb{P}\left[U\right]}\leq\infty$ for all $a\in\mathcal{A},x\in\mathcal{X}$.
We can show the existence using a similar discussion to 3.1. The structural function is obtained by the partial average over the bridge function as $\mathbb{E}\left[h(\tilde{a},X,W)\right]=f_{\mathrm{struct}}(\tilde{a})$ if we further assume either of the following:
Assumption A.4 ( 24, Assumption 2(iii)).
For test point $\tilde{a}\in\mathcal{A}$, we have $W\mathop{\perp\!\!\!\perp}A|X,Y^{(\tilde{a})}$.
Assumption A.5.
For any square integrable function $l:\mathcal{U}\to\mathbb{R}$, the following conditions hold for test point $\tilde{a}\in\mathcal{A}$ and for any $x\in\mathcal{X}$:
$\mathbb{E}\left[l(U)\mid A=\tilde{a},Y=y,X=x\right]=0\quad\forall y\in\mathcal{Y}\quad\Leftrightarrow\quad l(u)=0\quad\mathbb{P}\left[U\right]\text{-}\mathrm{a.e.}$
We may use either two-stage regression or the maximum moment restriction approach to obtain the bridge function. We state the closed-form solution for each approach.
Theorem A.6.
Given stage 1 samples $\{w_{i},a_{i},y_{i},x_{i}\}_{i=1}^{n}$ and stage 2 samples $\{\dot{a}_{i},\dot{y}_{i},\dot{x}_{i}\}_{i=1}^{m}$, and regularizing parameter $(\lambda,\eta)$, the solution of two-stage regression approach is given as
$\hat{h}(a,w,x)=\bm{\alpha}^{\top}\bm{k}(a,w,x),\quad\bm{\alpha}=(M+m\eta I)^{-1}\dot{\bm{y}}$
where $\dot{\bm{y}}=(\dot{y}_{1},\dots,\dot{y}_{m})^{\top}\in\mathbb{R}^{m}$, and $M\in\mathbb{R}^{m\times m}$ and $\bm{k}(a,w)\in\mathbb{R}^{m}$ are defined as
$M=K_{\dot{A}\dot{A}}\odot(B^{\top}K_{WW}B)\odot K_{\dot{X}\dot{X}},$
$\bm{k}(a,w,x)=\bm{k}_{\dot{A}}(a)\odot(B^{\top}\bm{k}_{W}(w))\odot\bm{k}_{\dot{X}}(x)$
$B=(K_{AA}\odot K_{YY}\odot K_{XX}+n\lambda I){}^{-1}(K_{A{\dot{A}}}\odot K_{Y\dot{Y}}\odot K_{X\dot{X}}).$
Here, for $F\in\{A,W,Y,X\}$, we denote the stage 1 kernel matrix as $K_{FF}=(k_{\mathcal{F}}(f_{i},f_{j}))_{ij}\in\mathbb{R}^{n\times n}$ where $k_{\mathcal{F}}$ and $f_{i}$ are the corresponding space and stage 1 samples. Similarly, for $\dot{F}\in\{\dot{A},\dot{Y},\dot{X}\}$, we denote the stage 2 kernel matrix as $K_{\dot{F}\dot{F}}=(k_{\mathcal{F}}(\dot{f}_{i},\dot{f}_{j}))_{ij}\in\mathbb{R}^{n\times n}$ where $k_{\mathcal{F}}$ and $f_{i}$ are the corresponding space and stage 2 samples. We further denote $K_{F\dot{F}}=(k_{\mathcal{F}}(f_{i},\dot{f}_{j}))_{ij}\in\mathbb{R}^{n\times m}$ and $\bm{k}_{\dot{A}}(a)=(k(\dot{a}_{i},a))_{i}\in\mathbb{R}^{m},\bm{k}_{W}(w)=(k(w_{i},w))_{i}\in\mathbb{R}^{n},\bm{k}_{\dot{X}}(s)=(k(\dot{x}_{i},x))_{i}\in\mathbb{R}^{m}$
Theorem A.7.
Given samples $\{w_{i},a_{i},y_{i},x_{i}\}_{i=1}^{n}$, the solution of maximum moment restriction approach is given as $\hat{h}(a,w)=\bm{\alpha}^{\top}\bm{k}(a,w)$, where weight $\bm{\alpha}\in\mathbb{R}^{n}$ is given as
$\bm{\alpha}=\sqrt{G}\left(\sqrt{G}L\sqrt{G}+n^{2}\eta I\right)^{-1}\sqrt{G}\bm{y}.$
Here, $\bm{y}=(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$ and
$L=K_{AA}\odot K_{WW}\odot K_{XX},\quad G=K_{AA}\odot K_{YY}\odot K_{XX},$
$\bm{k}(a,w)=(k_{\mathcal{A}}(a_{i},a)k_{\mathcal{W}}(w_{i},w)k_{\mathcal{X}}(x_{i},x))_{i}\in\mathbb{R}^{n}$
with $K_{FF}=(k_{\mathcal{F}}(f_{i},f_{j}))_{ij}\in\mathbb{R}^{n\times n}$ for $F\in\{W,A,Y,X\}$ and corresponding kernel function $k_{\mathcal{F}}$ and samples $f_{i}$, and $\sqrt{G}$ denotes the square root of $G=\sqrt{G}\sqrt{G}$.
The consistency of each estimation can be proved by a similar discussion presented in 4.5 and 4.7.
Appendix B Connection to PCL
In Proxy Causal Learning (PCL), we assume access to a treatment-inducing proxy variable $Z$, and an outcome-inducing proxy variable $W$, which satisfy the following structural assumption and completeness assumption.
Assumption B.1 (Structural Assumption [4, 13]).
We assume $Y\mathop{\perp\!\!\!\perp}Z|A,U$, and $W\mathop{\perp\!\!\!\perp}(A,Z)|U$.
Assumption B.2 (Completeness Assumption on Confounder [4, 13]).
Let $l:\mathcal{U}\to\mathbb{R}$ be any square integrable function $\|l\|_{\mathbb{P}\left[U\right]}<\infty$. The following conditions hold for all $a\in\mathcal{A}$
$\mathbb{E}\left[l(U)\mid A=a,W=w\right]=0\quad\forall w\in\mathcal{W}\quad\Leftrightarrow\quad l(u)=0\quad\mathbb{P}\left[U\right]\text{-}\mathrm{a.e.}$
$\mathbb{E}\left[l(U)\mid A=a,Z=z\right]=0\quad\forall z\in\mathcal{Z}\quad\Leftrightarrow\quad l(u)=0\quad\mathbb{P}\left[U\right]\text{-}\mathrm{a.e.}$
Here, $A,Y,U$ are the treatment, the outcome, and the unobserved confounder, respectively. We show that given Assumptions 4.3, 2.2, 2.5 and 2.3, these PCL assumptions hold with $Z=Y$.
Since $Y$ is deterministic given $A,U$ (Assumption 2.3), we have
$Y\mathop{\perp\!\!\!\perp}Y|A,U$
since the deterministic variables are independent of the deterministic variables by definition. Also, from Assumption 2.1, we have $W\mathop{\perp\!\!\!\perp}(Z,Y)|U$. Therefore, we have Assumption B.1. Furthermore, the first line of Assumption B.2 is equivalent to Assumption 2.2, and the second line of Assumption B.2 corresponds to Assumption 2.5 with $Z=Y$.
From this, we show that the outcome $Y$ can be used as the proxy $Z$ in PCL. We use this connection to derive the functional equation for bridge function (1) and its relation to structural function $f_{\mathrm{struct}}$. In PCL, the bridge function $h_{0}$ is defined as the solution of
$\mathbb{E}\left[h_{0}(a,W)|A,Z\right]=\mathbb{E}\left[Y|A,Z\right],$
and we have $f_{\mathrm{struct}}(a)=\mathbb{E}_{W}\left[h_{0}(a,W)\right]$. Now, if we replace the proxy $Z$ to the outcome $Y$, the right-hand side becomes just $Y$, and the left-hand side is $\mathbb{E}\left[h_{0}(a,W)|A,Y\right]$, which gives the definition in our bridge function (1). From this, we can derive the identification results (3.2) by using a similar discussion to the PCL identification.
Appendix C Use of Neural Network Feature
A number of works [25, 9, 7] proposed the deep learning methods for proxy causal learning, which can be applied to our settings by using the connection discussed in Appendix B. We expect these methods to achieve better empirical performances than kernel-based methods, especially when the data is high-dimensional as discussed in the original PCL setting [25, 9, 7]. We first discuss the two-stage regression approach using neural nets, and then introduce the neural version of the maximum moment restriction approach.
Two-stage Regression Approach
We can apply a similar method as DFPV [25] to derive the neural version of the two-stage regression approach. In DFPV, we model the bridge function $h_{0}(a,w)$ as
$h(a,w)=\bm{u}^{\top}(\bm{\psi}_{\theta_{A(2)}}(a)\otimes\bm{\psi}_{\theta_{W}}(w)),$
where $\bm{u}$ is weight parameters, $\otimes$ is the tensor product $\bm{a}\otimes\bm{b}=\mathrm{vec}(ab^{\top})$ and $\bm{\psi}_{\theta_{A(2)}},\bm{\psi}_{\theta_{W}}$ are neural nets parametrized by $\theta_{A(2)},\theta_{W}$, respectively. We minimize the loss $\mathcal{L}^{\mathrm{2SR}}$ to learn these parameters.
$\mathcal{L}^{\mathrm{2SR}}(h)=\mathbb{E}_{AY}\left[(Y-\bm{u}^{\top}(\bm{\psi}_{\theta_{A(2)}}(a)\otimes\mathbb{E}\left[\bm{\psi}_{\theta_{W}}(W)|A,Y\right]))\right]$
Similar to the SKPV, we need to model the conditional feature mean $\mathbb{E}\left[\bm{\psi}_{\theta_{W}}(W)|A,Y\right]$ to minimize the loss $\mathcal{L}^{\mathrm{2SR}}$. We model this using the different neural net feature maps.
$\mathbb{E}_{W|A=a,Y=y}\left[\bm{\psi}_{\theta_{W}}(W)\right]=\bm{V}(\bm{\phi}_{\theta_{A(1)}}(a)\otimes\bm{\phi}_{\theta_{Y}}(y)),$
where $\bm{V}$ are parameters, and $\bm{\phi}_{\theta_{A(1)}},\bm{\phi}_{\theta_{Z}}$ are neural nets parametrized by $\theta_{A(1)},\theta_{Z}$, respectively. Note we may use different neural nets in the treatment features $\bm{\phi}_{\theta_{A(1)}}$ and $\bm{\psi}_{\theta_{A(2)}}$.
As in SKPV, we learn $\mathbb{E}_{W|a,y}\left[\bm{\psi}_{\theta_{W}}(W)\right]$ in stage 1 regression and $h_{0}(a,w)$ in stage 2 regression, but in addition to the weights $\bm{u}$ and $\bm{V}$, we also learn the parameters of the feature maps.
Specifically, in stage 1, we learn $\bm{V}$ and parameters $\theta_{A(1)},\theta_{Z}$ by minimizing the following empirical loss:
^${\hat{\mathcal{L}}}^{\mathrm{cond}}(\bm{V},\theta_{A(1)},\theta_{Z})=\frac{1}{m}\sum_{i=1}^{m}\left\|\bm{\psi}_{\theta_{W}}(w_{i})-\bm{V}\left(\bm{\phi}_{\theta_{A(1)}}(a_{i})\otimes\bm{\phi}_{\theta_{Y}}(y_{i})\right)\right\|^{2}+\lambda\|\bm{V}\|^{2}.$
Although ^${\hat{\mathcal{L}}}^{\mathrm{cond}}$ depends on $\theta_{W}$, at this stage, we do not optimize $\theta_{W}$ with ^${\hat{\mathcal{L}}}^{\mathrm{cond}}$ as ${\bm{\psi}_{\theta_{W}}(w)}$ is the “target variable” in stage 1 regression. Given the minimizers $(\hat{\bm{V}},\hat{\theta}_{A(1)},\hat{\theta}_{Z})=\operatorname*{arg\,min}\hat{\mathcal{L}}_{1}$, we learn weights $\bm{u}$ and parameters $\theta_{W},\theta_{A(2)}$ by minimizing the empirical stage 2 loss,
^${\hat{\mathcal{L}}}^{\mathrm{2SR}}(\bm{u},\theta_{W},\theta_{A(2)})=\frac{1}{n}\sum_{i=1}^{n}\left(\dot{y}_{i}-\bm{u}^{\top}\left(\bm{\psi}_{\theta_{A(2)}}(\dot{a}_{i})\otimes\hat{\bm{V}}\left(\bm{\phi}_{\hat{\theta}_{A(1)}}(\dot{a}_{i})\otimes\bm{\phi}_{\hat{\theta}_{Y}}(\dot{y}_{i})\right)\right)\right)^{2}+\eta\|\bm{u}\|^{2}.$
Although the expression of ^${\hat{\mathcal{L}}}^{\mathrm{2SR}}$ does not explicitly contain $\theta_{W}$, it implicitly depends on $\theta_{W}$ as $(\hat{\bm{V}},\hat{\theta}_{A(1)},\hat{\theta}_{Z})$ is the solution of a minimization problem involving $\theta_{W}$. To deal with this implicit dependency, we may use the method proposed in Xu et al. [25], in which we ignore the dependency of $\theta_{W}$ on parameters $\hat{\theta}_{A(1)},\hat{\theta}_{Y}$, and compute the gradient via the closed-form solution of $\hat{V}$.
This gives the following learning procedure. First, we fix parameters in the adaptive feature maps $(\theta_{A(1)},\theta_{Y},\theta_{A(2)},\theta_{W})$. Given these parameters, optimal weights $\hat{\bm{V}},\hat{\bm{u}}$ can be learned by minimizing the empirical stage 1 loss ^${\hat{\mathcal{L}}}^{\mathrm{cond}}$ and empirical stage 2 loss ^${\hat{\mathcal{L}}}^{\mathrm{2SR}}$, respectively. These minimizations can be solved analytically, where the solutions are
$\hat{\bm{V}}(\bm{\theta})=\Psi_{1}^{\top}\Phi_{1}(\Phi_{1}^{\top}\Phi_{1}+m\lambda_{1}I)^{-1},$ $\hat{\bm{u}}(\bm{\theta})=\left(\Phi_{2}^{\top}\Phi_{2}+n\lambda_{2}I\right)^{-1}\Phi_{2}^{\top}\bm{y}_{2},$ (6)
where we denote $\bm{\theta}=(\theta_{A(1)},\theta_{Y},\theta_{A(2)},\theta_{W})$ and define matrices as follows:
$\Psi_{1}=\left[\bm{\psi}_{\theta_{W}}(w_{1}),\dots,\bm{\psi}_{\theta_{W}}(w_{m})\right]^{\top},$ $\Phi_{1}=[\bm{v}_{1}(a_{1},y_{1}),\dots,\bm{v}_{1}(a_{m},y_{m})]^{\top},$
$\bm{y}_{2}=[\dot{y}_{1},\dots,\dot{y}_{n}]^{\top},$ $\Phi_{2}=[\bm{v}_{2}(\dot{a}_{1},\dot{y}_{1}),\dots,\bm{v}_{2}(\dot{a}_{n},\dot{y}_{n})]^{\top},$
$\bm{v}_{1}(a,y)=\bm{\phi}_{\theta_{A(1)}}(a)\otimes\bm{\phi}_{\theta_{Y}}(y),$ $\bm{v}_{2}(a,y)=\bm{\psi}_{\theta_{A(2)}}(a)\otimes\left(\hat{\bm{V}}(\bm{\theta})\left(\bm{\phi}_{\theta_{A(1)}}(a)\otimes\bm{\phi}_{\theta_{Y}}(y)\right)\right).$
Given these weights $\hat{\bm{u}}(\bm{\theta}),\hat{\bm{V}}(\bm{\theta})$, we can update feature parameters by a gradient descent method with respect to the residuals of the loss of each stage, while regrading $\hat{\bm{V}}$ and $\hat{\bm{u}}$ as functions of parameters. Specifically, we take the gradient of the losses
^${\hat{\mathcal{L}}}^{\mathrm{cond}}(\bm{\theta})=\frac{1}{m}\sum_{i=1}^{m}\left\|\bm{\psi}_{\theta_{W}}(w_{i})-\hat{\bm{V}}(\bm{\theta})\left(\bm{\phi}_{\theta_{A(1)}}(a_{i})\otimes\bm{\phi}_{\theta_{Y}}(y_{i})\right)\right\|^{2}+\lambda\|\hat{\bm{V}}(\bm{\theta})\|^{2},$
^${\hat{\mathcal{L}}}^{\mathrm{2SR}}(\bm{\theta})=\frac{1}{n}\sum_{i=1}^{n}\left(\dot{y}_{i}-\hat{\bm{u}}(\bm{\theta})^{\top}\left(\bm{\psi}_{\theta_{A(2)}}(\dot{a}_{i})\otimes\hat{\bm{V}}(\bm{\theta})\left(\bm{\phi}_{\theta_{A(1)}}(\dot{a}_{i})\otimes\bm{\phi}_{\theta_{Y}}(\dot{y}_{i})\right)\right)\right)^{2}\!+\!\eta\|\hat{\bm{u}}(\bm{\theta})\|^{2},$
where $\hat{\bm{V}}(\bm{\theta}),\hat{\bm{u}}(\bm{\theta})$ are given in (6). Given these losses, $(\theta_{A(1)},\theta_{Y})$ are minimized with respect to ^${\hat{\mathcal{L}}}^{\mathrm{cond}}$, and $(\theta_{A(2)},\theta_{W})$ are minimized with respect to ^${\hat{\mathcal{L}}}^{\mathrm{2SR}}$. Xu et al. [25] reported that the learning procedure is stabilized by running several gradient descent steps on the stage 1 parameters $(\theta_{A(1)},\theta_{Y})$ before updating the stage 2 features $(\theta_{A(2)},\theta_{W})$.
Maximum Moment Restriction
In maximum moment restriction, we minimize the following loss to obtain the bridge function
$\hat{h}=\operatorname*{arg\,min}_{h}\max_{g}(\mathbb{E}\left[(Y-h(A,W))g(A,Y)\right])^{2}.$
In SPMMR, we assume both $h$ and $g$ to be in RKHS. There are two approaches to extend this to the neural network model. One is to model $h$ as a neural network while still assuming $g$ is in RKHS, which is considered in Kompa et al. [9]. The other is to model both $h,g$ as neural network functions as in Kallus et al. [7].
If we assume $g$ to be the RKHS function, we can learn the bridge function by minimizing the same loss $\mathcal{L}^{\mathrm{MMR}}$ to SPMMR.
$\hat{h}=\operatorname*{arg\,min}_{h}\mathbb{E}\left[\Delta_{Y,A,W}(h)\Delta_{Y^{\prime},A^{\prime},W^{\prime}}(h)k_{\mathcal{A}}(A,A^{\prime})k_{\mathcal{Y}}(Y,Y^{\prime})\right],$ (7)
where $\Delta_{Y,A,W}(h)=(Y-h(A,W))$ and $(W^{\prime},A^{\prime},Y^{\prime})$ are independent copies of $(W,A,Y)$, respectively. Empirically, we may add regularizer $\Omega$ and minimize the following loss.
$\hat{h}=\operatorname*{arg\,min}_{h}\frac{1}{n^{2}}\sum_{i,j=1}^{n}\Delta_{i}(h)\Delta_{j}(h)k_{\mathcal{A}}(a_{i},a_{j})k_{\mathcal{Y}}(y_{i},y_{j})+\eta\Omega(h)$
Kompa et al. [9] suggested to use the $L^{2}$-penalty on the parameter in $h$ as the regularizer.
When we model $g$ as a neural net function, we may still obtain the bridge function $h$ by solving (7). However, Kallus et al. [7] proposed the following loss, which is more computationally efficient.
$\hat{h}=\operatorname*{arg\,min}_{h}\max_{g}\frac{1}{n}\sum_{i=1}^{n}(y_{i}-h(a_{i},w_{i}))g(a_{i},y_{i})-\frac{\lambda}{n}\sum_{i=1}^{n}g^{2}(a_{i},y_{i})$
Note that the objective of minimax problem can be computed in $O(n)$ time, while the original loss (7) requires $O(n^{2})$. The second term is called stabilizer [7]. Kallus et al. [7] showed that stabilizers are different from regularizers, and generally should not let it vanish when the sample size grows.
Appendix D Details of Identifiability
In this appendix, we prove the propositions given in the main text. In the following, we assume that the spaces $\mathcal{U}$, $\mathcal{A}$, $\mathcal{W}$ are separable and completely metrizable topological spaces and equipped with Borel $\sigma$-algebras.
D.1 Proof of 3.3
Here, we present the proof of 3.3.
Proof.
Let $\mathcal{B}_{i}$ be the subset of $\{1,\dots,d_{\mathcal{U}}\}$ that
$\mathcal{B}_{i}=\{k\in\{1,\dots,d_{\mathcal{U}}\}~{}|~{}\mathbb{P}\left[U=k|a,Y=i\right]>0\}.$
Now, the $(i,j)$ element of matrix $P(Y|a,U)P(U|a,Y)$ is given as
$(P(Y|a,U)P(U|a,Y))_{ij}$ $=\sum_{k=1}^{n_{d_{\mathcal{U}}}}\mathbb{P}\left[Y=i|a,U=k\right]\mathbb{P}\left[U=k|a,Y=j\right]$
$=\sum_{k\in\mathcal{B}_{j}}\mathbb{P}\left[Y=i|a,U=k\right]\mathbb{P}\left[U=k|a,Y=j\right]$
$=\sum_{k\in\mathcal{B}_{j}}\mathbb{P}\left[Y=i|a,U=k\right]\frac{\mathbb{P}\left[Y=j|a,U=k\right]\mathbb{P}\left[U=k|a\right]}{\mathbb{P}\left[Y=j|a\right]}$
where the third equality holds from Bayes’ theorem. Since the outcome $Y$ is given deterministically ( Assumption 2.3), we have either $\mathbb{P}\left[Y=i|a,U=k\right]=1$ or $\mathbb{P}\left[Y=i|a,U=k\right]=0$, and from the definition of $\mathcal{B}_{j}$, we have $\mathbb{P}\left[Y=j|a,U=k\right]=1$. Since $\mathbb{P}\left[{Y=i|a,U=k}\right]=1$ only occurs for the single $i\in\{1,\dots,d_{\mathcal{Y}}\}$, for $i\neq j$, we have
$(P(Y|a,U)P(U|a,Y))_{ij}$ $=\sum_{k\in\mathcal{B}_{j}}\mathbb{P}\left[Y=i|a,U=k\right]\frac{\mathbb{P}\left[Y=j|a,U=k\right]\mathbb{P}\left[U=k|a\right]}{\mathbb{P}\left[Y=j|a\right]}$
$=0\quad(\because\mathbb{P}\left[Y=j|a,U=k\right]=1,\mathbb{P}\left[Y=i|a,U=k\right]=0).$
Furthermore, we have
$(P(Y|a,U)P(U|a,Y))_{ii}$ $=\sum_{k\in\mathcal{B}_{i}}\mathbb{P}\left[Y=i|a,U=k\right]\frac{\mathbb{P}\left[Y=i|a,U=k\right]\mathbb{P}\left[U=k|a\right]}{\mathbb{P}\left[Y=i|a\right]}$
$=\sum_{k\in\mathcal{B}_{i}}\frac{\mathbb{P}\left[Y=i|a,U=k\right]\mathbb{P}\left[U=k|a\right]}{\mathbb{P}\left[Y=i|a\right]}\quad(\because\mathbb{P}\left[Y=i|a,U=k\right]=1)$
$=\sum_{k\in\mathcal{B}_{i}}\mathbb{P}\left[U=k|a,Y=i\right]=1.$
Therefore, we have $P(Y|a,U)P(U|a,Y)=I_{d_{\mathcal{Y}}}$. ∎
D.2 Existence of the left inverse of $P(W|A=a,Y)$
Here, we discuss the condition of the existence of $H_{a}$, which is the left inverse of $P(W|A=a,Y)$. We show that $P(W|A=a,Y)$ is left invertible given the conditional independence assumption (Assumption 2.1), the deterministic confounding (3.3), and the discrete version of completeness assumption (Assumption 2.2) stated as follows.
Lemma D.1.
Under Assumption 2.2, for any $\bm{v}\in\mathcal{U}$,
$\bm{v}^{\top}P(U|A=a,W)=\bm{0}^{\top}\Leftrightarrow\bm{v}=\bm{0}$
holds for any $A=a$.
First, we show that
$P(W|A=a,Y)C_{1}=(P(Y|A=a,W)C_{2})^{\top}.$
where matrix $C_{1}$ and $C_{2}$ are defined as
$C_{1}$ $=\begin{bmatrix}\mathbb{P}\left[Y=1|A=a\right]&&&\\ &\mathbb{P}\left[Y=2|A=a\right]&&\\ &&\ddots&\\ &&&\mathbb{P}\left[Y=d_{\mathcal{Y}}|A=a\right]\end{bmatrix},$
$C_{2}$ $=\begin{bmatrix}\mathbb{P}\left[W=1|A=a\right]&&&\\ &\mathbb{P}\left[W=2|A=a\right]&&\\ &&\ddots&\\ &&&\mathbb{P}\left[W=d_{\mathcal{W}}|A=a\right]\end{bmatrix}$
Let $\bm{f}\in\mathbb{R}^{d_{\mathcal{W}}}$ and $\bm{g}\in\mathbb{R}^{d_{\mathcal{Y}}}$. Then, we have
$\bm{f}^{\top}$ $P(W|A=a,Y)C_{1}\bm{g}$
$=\sum_{i=1}^{d_{\mathcal{W}}}\sum_{j=1}^{d_{\mathcal{Y}}}f_{i}\mathbb{P}\left[W=i|A=a,Y=j\right]\mathbb{P}\left[Y=j|A=a\right]g_{j}$
$=\sum_{i=1}^{d_{\mathcal{W}}}\sum_{j=1}^{d_{\mathcal{Y}}}\sum_{k=1}^{d_{\mathcal{U}}}f_{i}\mathbb{P}\left[W=i|A=a,Y=j,U=k\right]\mathbb{P}\left[U=k|A=a,Y=j\right]\mathbb{P}\left[Y=j|A=a\right]g_{j}$
$=\sum_{i=1}^{d_{\mathcal{W}}}\sum_{j=1}^{d_{\mathcal{Y}}}\sum_{k=1}^{d_{\mathcal{U}}}f_{i}\mathbb{P}\left[W=i|A=a,U=k\right]\mathbb{P}\left[U=k|A=a,Y=j\right]\mathbb{P}\left[Y=j|A=a\right]g_{j}\quad(\because W\mathop{\perp\!\!\!\perp}Y|A,U)$
$=\sum_{i=1}^{d_{\mathcal{W}}}\sum_{j=1}^{d_{\mathcal{Y}}}\sum_{k=1}^{d_{\mathcal{U}}}f_{i}\mathbb{P}\left[W=i|A=a,U=k\right]\mathbb{P}\left[Y=j,U=k|A=a\right]g_{j}$
$=\sum_{i=1}^{d_{\mathcal{W}}}\sum_{j=1}^{d_{\mathcal{Y}}}\sum_{k=1}^{d_{\mathcal{U}}}f_{i}\mathbb{P}\left[W=i|A=a,U=k\right]\mathbb{P}\left[U=k|A=a\right]\mathbb{P}\left[Y=j|A=a,U=k\right]g_{j}$
Also,
$\bm{g}^{\top}$ $P(Y|A=a,W)C_{2}\bm{f}$
$=\sum_{i=1}^{d_{\mathcal{W}}}\sum_{j=1}^{d_{\mathcal{Y}}}f_{i}\mathbb{P}\left[W=i|A=a\right]\mathbb{P}\left[Y=j|A=a,W=i\right]g_{j}$
$=\sum_{i=1}^{d_{\mathcal{W}}}\sum_{j=1}^{d_{\mathcal{Y}}}\sum_{k=1}^{d_{\mathcal{U}}}f_{i}\mathbb{P}\left[W=i|A=a\right]\mathbb{P}\left[Y=j|A=a,W=i,U=k\right]\mathbb{P}\left[U=k|A=a,W=i\right]g_{j}$
$=\sum_{i=1}^{d_{\mathcal{W}}}\sum_{j=1}^{d_{\mathcal{Y}}}\sum_{k=1}^{d_{\mathcal{U}}}f_{i}\mathbb{P}\left[W=i,U=k|A=a\right]\mathbb{P}\left[Y=j|A=a,U=k\right]g_{j}\quad(\because W\mathop{\perp\!\!\!\perp}Y|A,U)$
$=\sum_{i=1}^{d_{\mathcal{W}}}\sum_{j=1}^{d_{\mathcal{Y}}}\sum_{k=1}^{d_{\mathcal{U}}}f_{i}\mathbb{P}\left[W=i|A=a,U=k\right]\mathbb{P}\left[U=k|A=a\right]\mathbb{P}\left[Y=j|A=a,U=k\right]g_{j}.$
From this, we can see that for all vector $\bm{f},\bm{g}$, we have
$\bm{f}^{\top}P(W|A=a,Y)C_{1}\bm{g}=\bm{g}^{\top}P(Y|A=a,W)C_{2}\bm{f},$
hence $P(W|A=a,Y)C_{1}=(P(Y|A=a,W)C_{2})^{\top}$. Now, in order to prove $P(W|A=a,Y)$ is left invertible, we need to show
$P(W|A=a,Y)\bm{\ell}=\bm{0}\Rightarrow\bm{\ell}=\bm{0}.$
Since $C_{1},C_{2}$ are invertible, we have
$P(W|A=a,Y)\bm{\ell}=\bm{0}$ $\Rightarrow P(W|A=a,Y)C_{1}C^{-1}_{1}\bm{\ell}=\bm{0}$
$\Rightarrow(C^{-1}_{1}\bm{\ell})^{\top}P(Y|A=a,W)C_{2}=\bm{0}^{\top}$
$\Rightarrow(C^{-1}_{1}\bm{\ell})^{\top}P(Y|A=a,U)P(U|A=a,W)=\bm{0}^{\top}$
$\Rightarrow((P(Y|A=a,U))^{\top}C^{-1}_{1}\bm{\ell})^{\top}P(U|A=a,W)=\bm{0}^{\top}$
$\Rightarrow(P(Y|A=a,U))^{\top}C^{-1}_{1}\bm{\ell}=\bm{0},$
where the last relation holds from D.1. From 3.3, we have
$(P(U|A=a,Y))^{\top}(P(Y|A=a,U))^{\top}=I_{d_{\mathcal{Y}}},$
and thus, we have
$(P(Y|A=a,U))^{\top}C^{-1}_{1}\bm{\ell}=\bm{0}$ $\Rightarrow(P(U|A=a,Y))^{\top}(P(Y|A=a,U))^{\top}C^{-1}_{1}\bm{\ell}=\bm{0}$
$\Rightarrow C^{-1}_{1}\bm{\ell}=\bm{0}$
$\Rightarrow\bm{\ell}=\bm{0}.$
This shows the existence of left inverse $H(a)$ under the condition of Assumption 2.1, 3.3 and D.1.
D.3 Existence of bridge function
First, we discuss conditions to guarantee the existence of the bridge function $h_{0}$. Let us consider the following operators:
$E_{a}:L^{2}(P_{W|A=a})\to L^{2}(P_{Y|A=a}),\ E_{a}f$ $\coloneqq\mathbb{E}\left[f(W)|A=a,Y=\cdot\right],$
$F_{a}:L^{2}(P_{Y|A=a})\to L^{2}(P_{W|A=a}),\ F_{a}g$ $\coloneqq\mathbb{E}\left[g(Y)|A=a,W=\cdot\right],$
Our goal is to show that $I(y)=y$ is in the range of $E_{a},$ i.e., we seek a solution to the inverse problem defined by
$E_{a}h=I(y).$ (8)
This suffices to prove the existence of the function $h,$ for if there exists a function $h^{*}_{a}$ for each $a\in{\cal A}$ such that
$\mathbb{E}\left[h^{*}_{a}|A=a,Y=\cdot\right]=y,$
we can define $h_{0}(a,w)\coloneqq h^{*}_{a}(w)$. We apply the discussion similar to Xu et al. [25, Appendix B], which ensures the existence using the following theorem.
Proposition D.2 (10, Theorem 15.18).
Let $\mathcal{X}$ and $\mathcal{Y}$ be Hilbert spaces. Let $E:\mathcal{X}\to\mathcal{Y}$ be a compact linear operator with singular system $\{(\mu_{n},\varphi_{n},g_{n})\}_{n=1}^{\infty}.$ The equation of the first kind
$E\varphi=f$
is solvable if and only if $f\in N(E^{*})^{\perp}$ and
$\sum_{n=1}^{\infty}\frac{1}{\mu_{n}^{2}}\lvert\langle f,g_{n}\rangle\rvert^{2}<\infty.$
Here, $N(E^{*})$ denotes the null space of the operator $E^{*}$. Then a solution is given by
$\phi=\sum_{n=1}^{\infty}\frac{1}{\mu_{n}}\langle f,g_{n}\rangle\varphi_{n}.$
To apply Proposition D.2, we make the following additional assumptions.
Assumption D.3.
For each $a\in{\cal A},$ the operator $E_{a}$ is compact with singular system $\{(\mu_{a,n},\varphi_{a,n},g_{a,n})\}_{n=1}^{\infty}.$
Assumption D.4.
The identity map $I(y)$ satisfies
$\sum_{n=1}^{\infty}\frac{1}{\mu_{a,n}^{2}}\lvert\langle I,g_{a,n}\rangle_{L^{2}(P_{Y|A=a})}\rvert^{2}<\infty,$
for a singular system $\{(\mu_{a,n},\phi_{a,n},g_{a,n})\}_{n=1}^{\infty}$ given in Assumption D.3.
Please refer to Xu et al. [25, Remark 1] for the discussion on these assumptions. It is easy to see that Assumptions D.3 and D.4 are required for using Proposition D.2. The remaining condition to show is that $I(y)$ is in $N(E_{a}^{*})^{\perp}.$ We show that the structural assumption (Assumption 2.1), completeness assumption (Assumption 2.2), and deterministic confounding (Assumption 2.3) imply the required condition.
Lemma D.5.
Under Assumptions 2.1, 2.2 and 2.3, the identity $I(y)$ is in the orthogonal complement of the null space $N(E_{a}^{*}).$
Proof.
We first show that the adjoint of $E_{a}$ is given by $F_{a}.$ For the operator $E_{a},$ any $f\in L_{2}(P_{W|A=a})$ and $g\in L_{2}(P_{Y|A=a}),$ we have
$\left\langle E_{a}f,g\right\rangle_{L_{2}(P_{Y|A=a})}$ $=\mathbb{E}_{Y|A=a}\left[\mathbb{E}\left[f(W)|A=a,Y\right]g(Y)\right]$
$=\mathbb{E}_{Y|A=a}\left[\mathbb{E}_{U|A=a,Y}\left[\mathbb{E}\left[f(W)|A=a,Y,U\right]\right]g(Y)\right]$
=$\overset{\mathrm{(a)}}{=}\mathbb{E}_{Y|A=a}\left[\mathbb{E}_{U|A=a,Y}\left[\mathbb{E}\left[f(W)|A=a,U\right]\right]g(Y)\right]$
$=\mathbb{E}_{U,Y|A=a}\left[\mathbb{E}\left[f(W)|A=a,U\right]g(Y)\right]$
$=\mathbb{E}_{U|A=a}\left[\mathbb{E}\left[f(W)|A=a,U\right]\mathbb{E}\left[g(Y)|A=a,U\right]\right]$
where (a) follows from $W\mathop{\perp\!\!\!\perp}(A,Y)|U$, which is from Assumption 2.1. Similarly,
$\left\langle f,F_{a}g\right\rangle_{L^{2}(P_{W|A=a})}=\mathbb{E}_{W|A=a}\left[f(W)\mathbb{E}\left[g(Y)|A=a,W\right]\right]$
$=\mathbb{E}_{W|A=a}\left[f(W)\mathbb{E}_{U|A=a,W}\left[\mathbb{E}\left[g(Y)|A=a,W,U\right]\right]\right]$
=$\overset{\mathrm{(b)}}{=}\mathbb{E}_{W|A=a}\left[f(W)\mathbb{E}_{U|A=a,W}\left[\mathbb{E}\left[g(Y)|A=a,U\right]\right]\right]$
$=\mathbb{E}_{W,U|A=a}\left[f(W)\mathbb{E}\left[g(Y)|A=a,U\right]\right]$
$=\mathbb{E}_{U|A=a}\left[\mathbb{E}\left[f(W)|A=a,U\right]\mathbb{E}\left[g(Y)|A=a,U\right]\right]$
$=\left\langle E_{a}f,g\right\rangle_{L_{2}(P_{Y|A=a})}.$
Again, (b) is given by $W\mathop{\perp\!\!\!\perp}A,Y|U$ from Assumption 2.1. For any $f^{*}\in N(E^{*}_{a})=N(F_{a}),$ by iterated expectations, we have
00 0 $=\mathbb{E}\left[f^{*}(Y)|A=a,W=\cdot\right]$
$=\mathbb{E}_{U}\left[\mathbb{E}\left[f^{*}(Y)|A,U,W\right]|A=a,W=\cdot\right]$
[…]
From Assumption 2.2,
$\mathbb{E}\left[l(U)\mid A=a,W=w\right]=0\quad\forall w\in\mathcal{W}\quad\Leftrightarrow\quad l(u)=0\quad P_{U}\text{-}\mathrm{a.e.}$
for all functions $l\in L^{2}(P_{U|A=a})$. Hence, (9) and Assumption 2.2 implies
$f^{*}(\gamma_{0}(a,U))=0\quad P_{U}\text{-}\mathrm{a.s.}$
Then, the inner product between $f^{*}$ and $I(y)$ is given as follows:
$\left\langle f^{*},I\right\rangle_{L^{2}(P_{Y|A=a})}$ $=\mathbb{E}_{Y|A=a}\left[f^{*}(Y)Y\right]$
$=\mathbb{E}_{U|A=a}\left[\mathbb{E}_{Y|U,A=a}\left[f^{*}(Y)Y\right]\right]$
=$\overset{(c)}{=}\mathbb{E}_{U|A=a}\left[f^{*}(\gamma_{0}(a,U))\gamma_{0}(a,U)\right]$
$=0.$
Here, (c) uses the fact that $Y$ is deterministic given $(a,u)$. Hence, we have
$I\in N(E_{a}^{*})^{\perp}.$
Now, we are able to apply Proposition D.2, leading to the 3.1.
Proof of 3.1.
By Lemma D.5, the identity $I(y)$ is in $N(E_{a}^{*})^{\perp}.$ Therefore, by Proposition D.2, under the given assumptions, there exists a solution to (8). Letting the solution be $h^{*}_{a}$ completes the proof. ∎
D.4 Identifiability in deterministic confounding
Here, we show that the bridge function $h_{0}$ can be used to compute the structural function.
Proof of 3.2.
First, we show that under Assumption 2.3, we have
$\mathbb{E}\left[\gamma_{0}(\tilde{a},U)|A=\tilde{a},Y=y\right]=y.$ (10)
Let $B(\tilde{a},y)=\{u\in\mathcal{U}|\gamma(\tilde{a},u)=y\}$. Then, we have
$\mathbb{E}\left[\gamma_{0}(\tilde{a},U)|A=\tilde{a},Y=y\right]$ $=\int\gamma_{0}(\tilde{a},U)\mathrm{d}P(U|A=\tilde{a},Y=y)$
=$=\int_{\neg B(\tilde{a},y)}\gamma_{0}(\tilde{a},U)\mathrm{d}P(U|A=\tilde{a},Y=y)+\int_{B(\tilde{a},y)}\gamma_{0}(a,U)\mathrm{d}P(U|A=\tilde{a},Y=y)$
=$=\int_{\neg B(\tilde{a},y)}\gamma_{0}(\tilde{a},U)\mathrm{d}P(U|A=\tilde{a},Y=y)+y\int_{B(\tilde{a},y)}\mathrm{d}P(U|A=\tilde{a},Y=y)$
Now, from Bayes’ theorem,
$P(U|A=\tilde{a},Y=y)\propto P(Y=y|U,A=\tilde{a})P(U|A=\tilde{a}).$
Since $Y$ is generated deterministically, the support of $P(Y=y|U,A=\tilde{a})$ is $B(\tilde{a},y)$. Therefore,
$\int_{\neg B(\tilde{a},y)}\gamma_{0}(\tilde{a},U)\mathrm{d}P(U|A=\tilde{a},Y=y)=0,\quad\int_{B(\tilde{a},y)}\mathrm{d}P(U|A=\tilde{a},Y=y)=1.$
Hence, we have (10). Using this, we have
$\mathbb{E}\left[h_{0}(\tilde{a},W)|A=\tilde{a},Y=y\right]-y=0$
[…]
If we assume Assumption 2.5, this means
$\mathbb{E}\left[h_{0}(\tilde{a},W)|U\right]=\gamma_{0}(\tilde{a},U),$
and thus
$f_{\mathrm{struct}}(\tilde{a})=\mathbb{E}_{U}\left[\gamma_{0}(\tilde{a},U)\right]=\mathbb{E}_{U}\left[\mathbb{E}\left[h_{0}(\tilde{a},W)|U\right]\right]=\mathbb{E}\left[h_{0}(\tilde{a},W)\right].$
We can draw a same conclusion if we instead assume Assumption 2.4 as
$\mathbb{E}\left[h(\tilde{a},W)\right]\overset{(a)}{=}\mathbb{E}_{Y^{(\tilde{a})}}\left[\mathbb{E}\left[h(\tilde{a},W)|Y^{(\tilde{a})},A=\tilde{a}\right]\right]\overset{(b)}{=}\int\mathbb{E}\left[h(\tilde{a},W)|Y=y,A=\tilde{a}\right]P(Y^{(\tilde{a})}=y)\mathrm{d}y\overset{(c)}{=}\mathbb{E}\left[Y^{(\tilde{a})}\right],$
where we used $A\mathop{\perp\!\!\!\perp}W|Y^{(\tilde{a})}$ in (a), “consisitency” $\mathbb{E}\left[h(\tilde{a},W)|Y^{(\tilde{a})}=y,A=\tilde{a}\right]=\mathbb{E}\left[h(\tilde{a},W)|Y=y,A=\tilde{a}\right]$ in (b), and $\mathbb{E}\left[h(\tilde{a},W)|A=\tilde{a},Y=y\right]=y$ in (c). ∎
D.5 Identifiability in stochastic counfounding
In this section, we derive the error bound in the stochastic confounding.
Proof.
The proof is similar to 3.2. First, we show that under Assumption 3.4, we have
$|\mathbb{E}\left[\gamma_{0}(\tilde{a},U)|A=\tilde{a},Y=y\right]-y|<M\quad\forall y\in\mathcal{Y}.$ (11)
Let $B(\tilde{a},y,\varepsilon)=\{u\in\mathcal{U}\mid|\gamma(\tilde{a},u)-y|\leq M\}$. Then, we have
$\mathbb{E}\left[\gamma_{0}(\tilde{a},U)|A=\tilde{a},Y=y\right]$ $=\int\gamma_{0}(\tilde{a},U)\mathrm{d}P(U|A=\tilde{a},Y=y)$
=$=\int_{\neg B(\tilde{a},y)}\gamma_{0}(\tilde{a},U)\mathrm{d}P(U|A=\tilde{a},Y=y)+\int_{B(\tilde{a},y)}\gamma_{0}(\tilde{a},U)\mathrm{d}P(U|A=\tilde{a},Y=y)$
Now, from Bayes’ theorem,
$P(U|A=\tilde{a},Y=y)\propto P(Y=y|U,A=\tilde{a})P(U|A=\tilde{a}).$
From Assumption 3.4, the support of $P(Y=y|U,A=\tilde{a})$ is $B(\tilde{a},y)$. Therefore,
$\int_{\neg B(\tilde{a},y)}\gamma_{0}(\tilde{a},U)\mathrm{d}P(U|A=\tilde{a},Y=y)=0,\quad\int_{B(\tilde{a},y)}\mathrm{d}P(U|A=\tilde{a},Y=y)=1$
Furthermore, we have
$\int_{B(\tilde{a},y)}\gamma_{0}(\tilde{a},U)\mathrm{d}P(U|A=\tilde{a},Y=y)$ $\geq\int_{B(\tilde{a},y)}(y-M)\mathrm{d}P(U|A=\tilde{a},Y=y)$
$=y-M.$
Similarly, we can also show
$\int_{B(\tilde{a},y)}\gamma_{0}(\tilde{a},U)\mathrm{d}P(U|A=\tilde{a},Y=y)\leq y+M.$
Hence, we have (11). Using this, we have
$|\mathbb{E}\left[h_{0}(\tilde{a},W)|A=\tilde{a},Y=y\right]-y|\leq M$
[…]
Hence, if we solve equation (1), we have
$f_{\mathrm{struct}}(\tilde{a})$ $=\mathbb{E}_{U}\left[\gamma_{0}(\tilde{a},U)\right]$
$\leq\mathbb{E}_{U}\left[\mathbb{E}\left[h_{0}(\tilde{a},W)|U\right]+M\Xi\right]$
$=\mathbb{E}_{W}\left[h_{0}(\tilde{a},W)\right]+M\Xi.$
Similarly, we have $f_{\mathrm{struct}}(\tilde{a})\geq\mathbb{E}_{W}\left[h_{0}(\tilde{a},W)\right]-M\Xi$. ∎
Appendix E Derivation
Two-stage Regression
Here, we present the proof of 4.1. First, we recall the empirical estimate of conditional mean embedding as follows.
Proposition E.1 ([21, Theorem 5]).
Given stage 1 samples $\{w_{i},a_{i},y_{i}\}$, the solution of stage 1 regression is given as
$\hat{\mu}_{W|A,Y}(a,y)=\sum_{i=1}^{n}\beta_{i}(a,y)\phi(w_{i}),\quad\bm{\beta}(a,y)=(K_{AA}\otimes K_{YY}+n\lambda I)^{-1}(\bm{k}_{A}(a)\odot\bm{k}_{Y}(y))$
where $\bm{k}_{\mathcal{A}}(a)=(k(a_{i},a))_{i}\in\mathbb{R}^{n},\bm{k}_{\mathcal{Y}}(y)=(k(y_{i},y))_{i}$.
Given this, we consider stage 2 regression
$\hat{h}=\operatorname*{arg\,min}_{h\in\mathcal{H}_{\mathcal{A}\mathcal{W}}}\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\left\langle h,\phi_{\mathcal{A}}(a_{i})\otimes\hat{\mu}_{W|A,Y}(a_{i},y_{i})\right\rangle_{\mathcal{H}_{\mathcal{A}\mathcal{W}}}\right)^{2}+\eta\|h\|^{2}_{\mathcal{H}_{\mathcal{A}\mathcal{W}}}$
can be seen as the kernel ridge regression with the kernel function $\tilde{k}$ is given as
$\tilde{k}((a,y),(a^{\prime},y^{\prime}))$ $=\left\langle\phi_{\mathcal{A}}(a)\otimes\hat{\mu}_{W|A,Y}(a,y),\phi_{\mathcal{A}}(a^{\prime})\otimes\hat{\mu}_{W|A,Y}(a^{\prime},y^{\prime})\right\rangle_{\mathcal{H}_{\mathcal{A}\mathcal{W}}}$
$=k_{\mathcal{A}}(a,a^{\prime})\left(\bm{\beta}^{\top}(a,y)K_{WW}\bm{\beta}(a^{\prime},y^{\prime})\right).$
Using the closed-form solution of kernel ridge regression (See [15, Chapter 14.4.3]) yields 4.1.
Equivalence to Mastouri and Zhu et al. [13]
Mastouri and Zhu et al. [13] considered a different way of deriving the closed-form of the two-stage regression.
Proposition E.2 (Mastouri and Zhu et al. [13, Proposition 2]).
Given stage 1 samples $\{w_{i},a_{i},y_{i}\}_{i=1}^{n}$ and stage 2 samples $\{\dot{a}_{i},\dot{y}_{i}\}_{i=1}^{m}$, and regularizing parameter $(\lambda,\eta)$, we have
$\hat{h}(a,w)=\bm{k}_{W}^{\top}(w)\Gamma\bm{k}_{\dot{A}}(a),\quad\Gamma\in\mathbb{R}^{n\times m},\quad\mathrm{vec}(\Gamma)=(B\bar{\otimes}I)(M+m\eta I)^{-1}\dot{\bm{y}}\in\mathbb{R}^{nm}$
where $\bar{\otimes}$ is the tensor product of associated columns of matrices with the same number of columns, and $\mathrm{vec}$ is row-wise vectorization as $\mathrm{vec}(\begin{bmatrix}a&b\\ c&d\end{bmatrix})=[a,b,c,d]^{\top}$. The definitions of $\bm{k}_{W},\bm{k}_{\dot{A}},B,M$ can be found in 4.1.
Although this learns $mn$ parameters, we show this solution is equivalent to 4.1 restated as follows.
$\hat{h}(a,w)=\bm{\alpha}^{\top}(\bm{k}_{\dot{A}}(a)\odot B^{\top}\bm{k}_{W}(w)),\quad\bm{\alpha}=(M+m\eta I)^{-1}\dot{\bm{y}}\in\mathbb{R}^{m}$
From the definition of $\bar{\otimes}$ and $\bm{\alpha}$, we have
$\mathrm{vec}(\Gamma)=\begin{bmatrix}B_{11}&&&\\ &B_{12}&&\\ &&\ddots&\\ &&&B_{1n}\\ B_{21}&&&\\ &B_{22}&&\\ &&\ddots&\\ &&&B_{2n}\\ &&\vdots&\\ B_{m1}&&&\\ &B_{m2}&&\\ &&\ddots&\\ &&&B_{mn}\\ \end{bmatrix}\bm{\alpha}.$
Hence, from definition of $\mathrm{vec}$, we have
$\Gamma=\begin{bmatrix}(\bm{B}_{1\cdot}\odot\bm{\alpha})^{\top}\\ (\bm{B}_{2\cdot}\odot\bm{\alpha})^{\top}\\ \vdots\\ (\bm{B}_{n\cdot}\odot\bm{\alpha})^{\top}\end{bmatrix}\in\mathbb{R}^{n\times m},$
where $\bm{B}_{i\cdot}$ is the $i$-th row vector of matrix $B$. Therefore, E.2 can be written as
$\hat{h}(a,w)=\sum_{i=1}^{m}k_{W}(w_{i},w)(\bm{B}_{i\cdot}\odot\bm{\alpha})^{\top}\bm{k}_{\dot{A}}.$
From linearity, we have
$\hat{h}(a,w)$ $=\sum_{i=1}^{m}k_{W}(w_{i},w)(\bm{B}_{i\cdot}\odot\bm{\alpha})^{\top}\bm{k}_{\dot{A}}$
$=\left(\sum_{i=1}^{m}\bm{B}_{i\cdot}k_{W}(w_{i},w)\odot\bm{\alpha}\right)^{\top}\bm{k}_{\dot{A}}$
$=(B^{\top}\bm{k}_{W}(w)\odot\bm{\alpha})^{\top}\bm{k}_{\dot{A}}$
$=(B^{\top}\bm{k}_{W}(w)\odot\bm{k}_{\dot{A}})^{\top}\bm{\alpha}$
The last equality holds for all vectors $(\bm{a}\odot\bm{b})^{\top}\bm{c}=(\bm{a}\odot\bm{c})^{\top}\bm{b}=\sum_{i}a_{i}b_{i}c_{i}$. Hence, we showed E.2 is equivalent to 4.1.
Maximum Moment Restriction
Here, we present the proof of 4.2. Let $\bm{h}\in\mathbb{R}^{n}=(h(a_{i},w_{i}))_{i}\in\mathbb{R}^{n}$. Then, we can rewrite empirical loss ^${\hat{\mathcal{L}}}^{\mathrm{MMR}}$ as
$\hat{h}$ $=\operatorname*{arg\,min}_{h}\frac{1}{n^{2}}\bm{h}^{\top}G\bm{h}-\frac{2}{n^{2}}\bm{y}^{\top}G\bm{h}+\frac{1}{n^{2}}\bm{y}^{\top}G\bm{y}+\eta\|h\|_{\mathcal{H}_{\mathcal{A}\mathcal{W}}}$
$=\operatorname*{arg\,min}_{h}\frac{1}{n}(\sqrt{G/n}\bm{y}-\sqrt{G/n}\bm{h})^{\top}(\sqrt{G/n}\bm{y}-\sqrt{G/n}\bm{h})+\eta\|h\|_{\mathcal{H}_{\mathcal{A}\mathcal{W}}}$
$=\operatorname*{arg\,min}_{h}\frac{1}{n}\sum_{i=1}^{n}((\sqrt{G/n}\bm{y})_{i}-(\sqrt{G/n}\bm{h})_{i})^{2}+\eta\|h\|_{\mathcal{H}_{\mathcal{A}\mathcal{W}}}.$
Here, we denote $(\bm{x})_{i}$ as the $i$-th element of vector $\bm{x}$. From reproducing characteristic, we have
$(\sqrt{G/n}\bm{h})_{i}=\left\langle\tilde{\phi}_{i},h\right\rangle,\quad\tilde{\phi}_{i}=\sum_{j=1}^{n}(\sqrt{G/n})_{ij}\phi_{\mathcal{A}\mathcal{W}}(a_{j},w_{j}),$
where $(\sqrt{G/n})_{ij}$ is the $(i,j)$-element of $\sqrt{G/n}$. Hence, we can regard this as kernel ridge regression, where the target is $(\sqrt{G/n}\bm{y})_{i}$ and the feature is $\tilde{\phi}_{i}$. Since
$\left\langle\tilde{\phi}_{i},\tilde{\phi}_{j}\right\rangle_{\mathcal{H}_{\mathcal{A}\mathcal{W}}}=(\sqrt{G/n}L\sqrt{G/n})_{ij},$
we have the closed-form solution as $\hat{h}=\bm{\beta}^{\top}\tilde{\bm{k}}(a,w)$
$\bm{\beta}$ $=(\sqrt{G/n}L\sqrt{G/n}+n\eta I)^{-1}\sqrt{G/n}\bm{y}$
$=\sqrt{n}(\sqrt{G}L\sqrt{G}+n^{2}\eta I)^{-1}\sqrt{G}\bm{y}$
$\tilde{\bm{k}}(a,w)$ $=\begin{bmatrix}\left\langle\tilde{\psi}_{1},\phi_{\mathcal{A}\mathcal{W}}(a,w)\right\rangle_{\mathcal{H}_{\mathcal{A}\mathcal{W}}}\\ \left\langle\tilde{\psi}_{2},\phi_{\mathcal{A}\mathcal{W}}(a,w)\right\rangle_{\mathcal{H}_{\mathcal{A}\mathcal{W}}}\\ \vdots\\ \left\langle\tilde{\psi}_{n},\phi_{\mathcal{A}\mathcal{W}}(a,w)\right\rangle_{\mathcal{H}_{\mathcal{A}\mathcal{W}}}\end{bmatrix}$
$=\sqrt{G/n}\bm{k}(a,w)$
This yields the closed-form solution in 4.2.